### v2

# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

Abstract
We introduce a new instrumental variable (IV) estimator for heterogeneous treatment effects in the presence of endogeneity. Our estimator is based on double/debiased machine learning (DML) and uses efficient machine learning instruments (MLIV) and kernel smoothing. We prove consistency and asymptotic normality of our estimator and also construct confidence sets that are more robust towards weak IV. Along the way, we also provide an accessible discussion of the corresponding estimator for the homogeneous treatment effect with efficient machine learning instruments. The methods are evaluated on synthetic and real datasets and an implementation is made available in the R package IVDML.
1 Introduction
A large part of research in causal inference and econometrics deals with estimation and inference for treatment effects based on observational data. As an example, suppose that we want to understand the effect of years of education on wages, as for example investigated by [10]. In this case, a randomized controlled trial is infeasible, since one cannot randomly assign years of education to students and one needs to rely on observational data alone. With purely observational data, a major issue is the potential endogeneity of treatment and response. For example, if there are unmeasured confounding factors that influence both the years of education as well as the wage (for example what can be summarized by “ability”), there will be endogeneity. A second issue that often appears is heterogeneity of the treatment effect. For example, the effect of education on wage could be different between men and women, in different geographical locations or it could depend on the age or the experience of the worker.
1.1 Our Contribution
We provide a contribution that deals with both of these issues and propose a novel procedure for estimation and inference of a heterogeneous treatment effect in the presence of endogeneity using an instrumental variable. Concretely, we consider the model
$Y_{i}=\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i},\,i=1,\ldots,N$ (1)
for a response variable $Y_{i}$, a univariate continuous covariate $V_{i}$, the treatment variable $D_{i}$, covariates $X_{i}$ and an error term $\epsilon_{i}$. We focus here on univariate $V_{i}$, but the extension to the multivariate case is conceptually straightforward, see also the discussion in Section 5. The goal is to perform inference for $\beta(v)$ for some specific value $v$. The error $\epsilon_{i}$ satisfies $\mathbb{E}[\epsilon_{i}|X_{i}]=0$ but we have endogeneity, meaning that $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]\neq 0$. However, we assume that we have access to an instrumental variable $Z_{i}$ that satisfies $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$. Our method is based on double/debiased machine learning [16] and uses a kernel smoothing approach to model the heterogeneity in the continuous variable $V_{i}$. We give conditions under which our method yields consistency and asymptotically valid confidence sets. Our main focus is neither to provide generic meta-algorithms for heterogenous treatment effect estimation in the spirit of [33, 25, 44] nor a detailed investigation of optimality properties [29, 30]. We rather want to introduce a simple, interpretable and easy-to-use method that works with user-chosen machine learning for nuisance function estimation. Nevertheless, many extensions of our method are possible by replacing kernel smoothing with other nonparametric regression or machine learning methods. Along the way, we also provide a discussion of some aspects of double machine learning with instrumental variables for homogeneous treatment effects (i.e., $\beta(\cdot)\equiv\beta$ is a constant) which, to our knowledge, have not yet been comprehensively discussed. Concretely, we discuss the usage of machine learning instruments and weak-instrument-robust confidence intervals. Finally, we provide an implementation of our method in the R package IVDML, which is available on CRAN [38].
1.2 Partially Linear IV Models
If the treatment effect $\beta(\cdot)$ in (1) is constant (i.e., does not depend on $V_{i}$), the model reduces to a standard partially linear model with endogeneity (i.e. $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]\neq 0$)
$Y_{i}=\beta D_{i}+g(X_{i})+\epsilon_{i},\,i=1,\ldots,N,$ (2)
where we assume the availability of an instrument $Z_{i}$. In our context, an instrument $Z_{i}$ must satisfy $\mathbb{E}[D_{i}|Z_{i},X_{i}]\neq\mathbb{E}[D_{i}|X_{i}]$ and $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$. The first condition is the relevance condition (the instrument is associated with the treatment conditionally on the covariates), and the second condition is the validity condition (the instrument is not associated with unmeasured confounders and affects the outcome only through the treatment conditionally on the covariates). Classical two-stage least squares instrumental variables theory, as for example discussed in many textbooks on econometrics, e.g. [54, 42], assumes that $g(X_{i})$ is a linear function in $X_{i}$ and utilizes only linear versions of relevance and validity. The validity of an instrument can in general not be tested but must be argued using domain knowledge. As noted for example by [19, 11], when researchers argue for the validity of instruments they often implicitly argue for mean independence and not for uncorrelatedness, so our stronger mean independence assumption is not too restrictive. On the other hand, the relevance condition $\mathbb{E}[D_{i}|Z_{i},X_{i}]\neq\mathbb{E}[D_{i}|X_{i}]$ is weaker than the classical assumption of correlation between $D_{i}$ and $Z_{i}$.
There is a large body of work that considers the partially linear model (2) in the endogenous setting, see for example [34, 21] for a review. Double/debiased machine learning (DML) [16] provides a general framework for performing inference on low-dimensional parameters that are identified using moment restrictions in the presence of potentially high-dimensional nuisance functions. If one uses a score function that is Neyman-orthogonal and sample-splitting with cross-fitting, one can use estimators for the nonparametric part of the partially linear model that converge comparably slow, which opens the possibility to use many modern machine learning methods and potentially high-dimensional covariates. As one concrete application, [16] apply their methodology to the partially linear IV model (2). DML in the context of model (2) has been studied more comprehensively by [21], allowing for multivariate $D_{i}$ and $Z_{i}$, potential overidentification and using a regularization scheme to obtain better finite sample performance and smaller confidence intervals.
An omnipresent issue with IV estimators is that they have large variance if the instrument is not very strong. Under the mean independence assumption $\mathbb{E}[\epsilon_{i}|X_{i},Z_{i}]=0$, every transformation $\zeta(Z_{i},X_{i})$ of the instrument, and the covariates is also a valid instrument and one can search for the transformation that leads to the smallest variance of the estimator. This fits into the larger context of semiparametric efficiency, where one seeks the estimator with the smallest asymptotic variance [46, 45].
The semiparametrically efficient estimator in the context of model (2) is for example presented in [47]. While semiparametrically efficient estimators are appealing from a theoretical perspective, they are often not recommended in practice, since they are too complicated to be estimated well with finite samples, see for example [56] for a related discussion.
For this reason, one often resorts to estimators that only achieve the semiparametric efficiency under the homoscedasticity assumption $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const}$. In this case, the transformation leading to the smallest asymptotic variance of the estimator is $\zeta(Z_{i},X_{i})=\mathbb{E}[D_{i}|Z_{i},X_{i}]$, which means that one can learn the optimal instruments using machine learning [6, 11, 31, 26]. Somewhat surprisingly, the solutions for model (2) using DML, do not use these ideas but use the instruments linearly [16, 21]. To our knowledge, the only place in the literature where double machine learning is combined with these efficient instruments is in [2], which presents a STATA implementation of DML that has one particular function that uses these efficient instruments (see Section 3.2 there), but the asymptotic results are not discussed in detail.
We will use the same idea of learning the efficient instruments from data and using DML for our treatment of the heterogeneous treatment effect in model (1). Along the way, we will also provide some additional details for the homogeneous treatment effect in model (2) that have not yet been comprehensively discussed in the literature.
1.3 Varying Coefficient Models
Another type of models related to (1) are varying coefficient models, see, e.g., [18, 27, 22, 23]. These are linear models, where the coefficient vector can vary depending on an additional variable, i.e.,
$Y_{i}=\gamma(U_{i})^{T}W_{i}+\epsilon_{i}.$ (3)
Moreover, varying coefficient models have also been studied in the endogenous setting with instrumental variables [9, 8, 43]. Structurally, our model (3) is related to model (1) by setting $U_{i}=(V_{i},X_{i})$, $W_{i}=(D_{i},1)$ and $\gamma(\cdot)=(\beta(\cdot),g(\cdot))$. However, there are important differences: we are only interested in conducting inference for the single component function $\beta(\cdot)$, whereas $g(\cdot)$ is a (potentially high-dimensional) nuisance function that allows to include additional control variables in a flexible fashion with arbitrary machine learning. We are not interested in estimating or conducting inference for the function $g(\cdot)$. In contrast, with varying coefficient models, typically, all the component functions are of equal interest and are modeled using low-dimensional techniques. Hence, we will look at model (1) not through the lens of varying coefficient models but heterogeneous treatment effects instead.
1.4 Heterogeneous Treatment Effects
When the heterogeneous treatment effect $\beta(\cdot)$ is of interest, there is a large body of works, that consider (variants of) model (1) but assume no endogeneity, that is $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]=0$ and there is no need for an instrument. An overview can be found in Chapters 14 and 15 of [17]. Most works consider a binary treatment $D_{i}$ taking values in $\{0,1\}$. In that case, $\beta(V_{i})$ in (1), is the Conditional Average Treatment Effect (CATE) $\beta(V_{i})=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|V_{i}]$, where $Y_{i}(1)$ and $Y_{i}(0)$ are the potential outcomes of $Y_{i}$ that would be observed when the treatment is assigned to be $D_{i}=0$ or $D_{i}=1$, respectively [17, 37]. We restrict ourselves here to the literature which is most related to our contribution.
[…]
1.5 Outline
The rest of the paper is structured as follows. In Section 2, we focus on the homogeneous treatment effect with machine learning instruments. This section mainly serves as a preparation and motivation for the following discussion of the heterogeneous treatment effect, but we also emphasize some aspects that have not been discussed in detail until now, most notably the construction of weak-instrument robust confidence sets. In Section 3, we consider the heterogeneous treatment effect model and introduce our estimator. Moreover, we give conditions under which we can perform inference for the heterogeneous treatment effect. In Section 4, we perform a simulation study and apply our method to two well-known instrumental variable datasets. Proofs and some additional simulation results can be found in the appendix.
1.6 Notation
For two sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ of positive real numbers, we write $a_{n}\lesssim b_{n}$ if there exists a constant $C>0$ such that for all $n\in\mathbb{N}$, $a_{n}\leq Cb_{n}$. We write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. We write $a_{n}\ll b_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. Moreover, we use stochastic order symbols: for a sequence of random variables $(X_{n})_{n\in\mathbb{N}}$ and a sequence of positive real numbers $(a_{n})_{n\in\mathbb{N}}$, we write $X_{n}=o_{P}(a_{n})$ if for all $\epsilon>0$, $\mathbb{P}(|X_{n}|/a_{n}\geq\epsilon)\to 0$, as $n\to\infty$. We write $X_{n}=O_{P}(a_{n})$ if $\limsup_{n\to\infty}\mathbb{P}(|X_{n}|/a_{n}\geq M)\to 0$, as $M\to\infty$. Equalities involving conditional expectations should be understood to hold almost surely.
2 Homogeneous Treatment Effect
In this section, we review and add to the discussion of DML in the partially linear IV model. Concretely, we observe $N$ i.i.d. realizations of $(Y_{i},D_{i},X_{i},Z_{i})$ from the model
$Y_{i}=\beta D_{i}+g(X_{i})+\epsilon_{i}$ (4)
for random variables $Y_{i}\in\mathbb{R}$, $D_{i}\in\mathbb{R}$, $X_{i}\in\mathbb{R}^{p}$ and the (unobserved) random error $\epsilon_{i}\in\mathbb{R}$ that is correlated with $D_{i}$ making the treatment $D_{i}$ endogeneous (i.e. $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]\neq 0$). To deal with the endogeneity, we assume the availability of an instrumental variable $Z_{i}\in\mathbb{R}^{d}$ satisfying Assumption 1 below. We allow the data generating process to vary with $N$, even if it is not reflected in the notation. This is needed to consider settings where the strength of the instrument $Z_{i}$ decreases with $N$. Furthermore, it allows for $X_{i}$ to be high-dimensional, i.e. $p\to\infty$ as $n\to\infty$. We assume that the instrument $Z_{i}$ satisfies the following conditions.
Assumption 1.
  1. 1.
    $\mathbb{E}[D_{i}|Z_{i},X_{i}]\neq\mathbb{E}[D_{i}|X_{i}]$,
  2. 2.
    $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$.
The first item states that the instrument is associated with the treatment conditionally on the covariates. For the second item, note that in model (4), we can assume without loss of generality that $\mathbb{E}[\epsilon_{i}|X_{i}]=0$, since we can otherwise change $g(X_{i})\leftarrow g(X_{i})+\mathbb{E}[\epsilon_{i}|X_{i}]$. Then, assuming $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]$ means that $Z_{i}$ is a valid instrument (see also the discussion in Section 1.2).
If the dimension of $Z_{i}$ is $d=1$, [16] and [21] identify $\beta$ as
$\beta=\frac{\mathbb{E}[(Y_{i}-\mathbb{E}[Y_{i}|X_{i}])(Z_{i}-\mathbb{E}[Z_{i}|X_{i}])]}{\mathbb{E}[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(Z_{i}-\mathbb{E}[Z_{i}|X_{i}])]},$ (5)
provided that $\mathbb{E}[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(Z_{i}-\mathbb{E}[Z_{i}|X_{i}]]\neq 0$. To construct an estimator, one replaces the inner conditional expectations by machine learning estimates and the outer expectations by sample means using a cross-fitting scheme (see details below). With this, one obtains an asymptotically normal estimator with variance that can be consistently estimated (see below). However, since $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$, one can in principle replace $Z_{i}$ by $\zeta(Z_{i},X_{i})$ for an arbitrary function $\zeta$ of $Z_{i}$ and $X_{i}$, i.e.
$\beta=\frac{\mathbb{E}[(Y_{i}-\mathbb{E}[Y_{i}|X_{i}])(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])]}{\mathbb{E}[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])]}.$ (6)
Note that here, we do not have the restriction $d=1$ of univariate $Z_{i}$ anymore but the function $\zeta$ should be real-valued. A natural goal is to find a transformation $\zeta$ that is easy to obtain and such that the resulting estimator has a small variance. We use $\zeta(Z_{i},X_{i})=f(Z_{i},X_{i})\coloneqq\mathbb{E}[D_{i}|Z_{i},X_{i}]$, which can be easily obtained by nonlinear regression of $D_{i}$ versus $(Z_{i},X_{i})$ and coincides with the semiparametrically efficient instrument (see e.g. equation (9) in [47]) for the homoscedastic case $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const.}$ This means that under homoscedasticity, the estimator based on the identification (6) achieves minimum asymptotic variance with $\zeta=f$ (see also Section 2.5 below).
In the literature, this approach of learning so-called optimal instruments is popular [6, 11, 31, 26], see also section 9.2 in [48] for an accessible discussion. However, to our knowledge, the only place in the literature where it is applied to double machine learning is [2], where the software package ddml for DML in Stata is presented which includes a function for the flexible partially linear IV model which coincides with the approach outlined above. In the following, we present the corresponding theory in our notation with the aim of motivating and facilitating the discussion of the heterogeneous treatment effect model in Section 3.
2.1 Construction
We now present the construction of the estimator using the cross-fitting scheme. We assume that we observe $N$ i.i.d. observations $(Y_{i},D_{i},X_{i},Z_{i})_{i=1}^{N}$ of model (4). Let us define the quantities
$f(Z_{i},X_{i})$ $\coloneqq\mathbb{E}[D_{i}|Z_{i},X_{i}],$
$\phi(X_{i})$ $\coloneqq\mathbb{E}[D_{i}|X_{i}]=\mathbb{E}[f(Z_{i},X_{i})|X_{i}],$
$l(X_{i})$ $\coloneqq\mathbb{E}[Y_{i}|X_{i}].$
Fix $K\in\mathbb{N}$ (independent of $N$). Let $\left\{I_{k}\right\}_{k=1}^{K}$ be a $K$-fold random partition of the observation indices $N]=\{1,\ldots,N\}$ such that each fold has size $|I_{k}|=n=N/K$ (for simplicity, we assume that $N/K$ is an integer). For $k=1,\ldots,K$, let $\hat{f}^{k}$ and $\hat{l}^{k}$ be estimates of $f$ and $l$ and let $\hat{\phi}_{1}^{k}$ and $\hat{\phi}_{2}^{k}$ be estimates of $\phi$, where $\hat{f}^{k}$, $\hat{l}^{k}$, $\hat{\phi}_{1}^{k}$ and $\hat{\phi}_{2}^{k}$ are obtained by machine learning only using the data $I_{k}^{c}$ (i.e. on all data apart from the $k$th fold). The reason for having two different estimators $\hat{\phi}_{1}^{k}$ and $\hat{\phi}_{2}^{k}$ for $\phi$ will be made clear below in Remark 2.
Define the residuals
$R_{Y,i}^{k}$ $\coloneqq Y_{i}-\hat{l}^{k}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K$ (7)
$R_{D,i}^{k}$ $\coloneqq D_{i}-\hat{\phi}_{1}^{k}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K$ (8)
$R_{f,i}^{k}$ $\coloneqq\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K.$ (9)
and define the estimator
$\hat{\beta}=\frac{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{f,i}^{k}}{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}}.$ (10)
Essentially, this is a two-stage least squares estimator of $\left(R_{Y,i}^{k}\right)_{i,k}$ on $\left(R_{D,i}^{k}\right)_{i,k}$ with instrument $\left(R_{f,i}^{k}\right)_{i,k}$ combined with the cross-fitting aggregation. This estimator is the same as equation (14) in [2].
Remark 1.
This estimator is the cross-fitting sample version of (6) for $\zeta(Z_{i},X_{i})=f(Z_{i},X_{i})=\mathbb{E}[D_{i}|Z_{i},X_{i}]$. The estimator is the analogue of the “DML2”-type estimator from [16]. The construction and analysis of the “DML1”-type estimator would be completely analogous, however, we omit this here, as the “DML2”-type estimator is typically preferred as it is more stable.
Remark 2.
The estimators $\hat{f}^{k}$, $\hat{l}^{k}$ and $\hat{\phi}_{1}^{k}$ can be obtained by directly fitting a machine learning model of $D_{i}$ vs. $(Z_{i},X_{i})$, $Y_{i}$ vs. $X_{i}$ and $D_{i}$ vs. $X_{i}$, respectively, using only $i\in I_{k}^{c}$. In principle, one could use the same direct estimator $\hat{\phi}_{2}^{k}=\hat{\phi}_{1}^{k}$ in the definition of the residuals $R_{f,i}^{k}$. However, similarly to [2], we recommend to use a two-stage approach for $\hat{\phi}_{2}^{k}$: after obtaining $\hat{f}^{k}$ on $I_{k}^{c}$, regress $(\hat{f}^{k}(Z_{i},X_{i}))_{i\in I_{k}^{c}}$ on $(X_{i})_{i\in I_{k}^{c}}$ to obtain $\hat{\phi}^{k}_{2}$, which is a different estimate of $\phi$ than $\hat{\phi}_{1}^{k}$. The intuition behind using $\hat{\phi}_{k}^{2}$ instead of $\hat{\phi}_{k}^{1}$ in the definition (9) of $R_{f,i}^{k}$ is that due to misspecification it might happen that $\hat{f}^{k}$ does not converge to the true $f$ but to some function $\zeta$. Then, we are in the setting of the identification (6) and should subtract an estimate of $\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$ rather than an estimate of $\phi(X_{i})=\mathbb{E}[D_{i}|X_{i}]$. Hence, we expect the estimator $\hat{\beta}$ with $R_{f,i}^{k}$ based on $\hat{\phi}_{2}^{k}$ to be more robust against poor estimation of $f$.
2.2 Variance Estimator
In Section 2.4 below, we will see conditions under which the estimator $\hat{\beta}$ is asymptotically normal, $\sqrt{N}(\hat{\beta}-\beta)/\sigma\to\mathcal{N}(0,1)$ in distribution and that the asymptotic variance $\sigma^{2}$ can be consistently estimated by
$\hat{\sigma}^{2}=\frac{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\hat{\beta}R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}}{\left(\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}\right)^{2}}.$ (11)
Consequently, an asymptotically valid $(1-\alpha)$-confidence interval is given by
$\left[\hat{\beta}-z_{1-\alpha/2}\hat{\sigma}/\sqrt{N},\,\hat{\beta}-z_{1-\alpha/2}\hat{\sigma}/\sqrt{N}\right],$ (12)
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of a standard Gaussian distribution.
2.3 Robust Confidence Intervals
The confidence interval constructed in (12) is not very robust to weak instrumental variables. In our case, weak IV means that $f(Z_{i},X_{i})-\phi(X_{i})$ is close to zero, i.e. there is no strong effect of $Z_{i}$ on $D_{i}$ conditional on $X_{i}$. In such cases, it can be helpful to use an alternative construction of confidence intervals. This construction is similar to confidence intervals obtained by inverting the Anderson-Rubin test [3]. We apply the method that is described in Chapter 13.3 in [17] based on identification (5) to our estimator.
For $\beta\in\mathbb{R}$, define the functions
$\hat{Q}(\beta)=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\beta R_{D,i}^{k}\right)R_{f,i}^{k}$ (13)
and
$\widehat{SE}_{Q}(\beta)=\sqrt{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{n}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\beta R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}-\hat{Q}(\beta)^{2}}.$ (14)
Under suitable conditions (see Section 2.4 below), under the null hypothesis $H_{0}:\beta=\beta_{0}$ (i.e. if the data is generated according to model (4) with $\beta=\beta_{0}$), we have that
$\frac{\sqrt{N}\hat{Q}(\beta_{0})}{\widehat{SE}_{Q}(\beta_{0})}\to\mathcal{N}(0,1)\text{ in distribution (under $H_{0}:\beta=\beta_{0}$)}$
and consequently,
$\mathbb{P}\left(\beta_{0}\in\left\{\beta:|\hat{Q}(\beta)|\leq z_{1-\alpha/2}\widehat{SE}_{Q}(\beta)/\sqrt{N}\right\}\right)\to 1-\alpha,$
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of a standard Gaussian distribution. Hence, the set $\mathcal{C}_{N}^{\alpha}=\left\{\beta:|\hat{Q}(\beta)|\leq z_{1-\alpha/2}\widehat{SE}_{Q}(\beta)/\sqrt{N}\right\}$ is an asymptotic $(1-\alpha)$-confidence set for $\beta$. The set $\mathcal{C}_{N}^{\alpha}$ is defined as the sublevel set of a parabola in $\beta$. Hence it is possible to derive an explicit expression, see Appendix C.1 for the expression in the heterogeneous case (the homogeneous case is analogue). Moreover, one can show that under some assumptions on the strength of the IV, the confidence interval $\mathcal{C}_{N}^{\alpha}$ is asymptotically equivalent to the classical confidence interval given in (12), see Appendix C.2 for the derivation in the heterogeneous case (the homogeneous case is analogue). Hence, one does not lose asymptotic efficiency from using the more robust version and we would recommend to use the robust confidence sets in practice. However, from a purely theoretical point of view, the advantage is not large: we still need that the nuisance function estimates converge fast enough, and essentially, the conditions on the convergence of the nuisance function estimates get scaled by the IV strength, which means that with weak IV, we get stronger requirements on how fast the nuisance function estimates need to converge, see also Remark 3. Nevertheless, in simulations one can observe that the robust confidence intervals do not suffer from undercoverage in settings with weak IV where the coverage of the standard confidence intervals breaks down (see Appendix E.1).
2.4 Summary of Results for the Homogeneous Treatment Effect
In this section, we give a summary of the theoretical results that one can obtain for the homogeneous treatment effect. These results essentially follow from the general theory in [16], since the score
$\psi(Y,D,X,Z;\beta,l,f,\phi)=\left(Y-l(X)-\beta(D-\phi(X))\right)\left(f(Z,X)-\phi(X)\right)$
satisfies the Neyman orthogonality property [2], but they can also be proved by directly analyzing the expressions (10) and (11) by essentially following the proofs that we will give for the heterogeneous treatment effects with some slight modifications. Additionally to some conditions on the moments of the random variables, the main requirements are that the nuisance functions $l$, $\phi$ and $f$ can be estimated at a sufficiently fast rate. Concretely, define
$\sigma_{Q}^{2}=\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}],$ (15)
which can be seen as a measure of IV strength on the population level. Note that $\sigma_{Q}^{2}$ is unnormalized, i.e. we do not divide by the noise level $\mathbb{E}[\epsilon_{i}^{2}]$. In contrast to the standard theory from [16] (Assumption 3.4. (d) there), we do not assume that $\sigma_{Q}^{2}$ is bounded away from 0, but we have to rescale the conditions on the convergence rates. We write $\|f-\hat{f}^{k}\|_{L_{2}}$ for the $L_{2}$-norm of $f-\hat{f}^{k}$. Note that $f$ is fixed and $\hat{f}^{k}$ depends on the data in $I_{k}^{c}$. Hence, $\|f-\hat{f}^{k}\|_{L_{2}}$ is random. More precisely, it is given by
$\|f-\hat{f}^{k}\|_{L_{2}}^{2}=\mathbb{E}\left[(f(Z,X)-\hat{f}^{k}(Z,X))^{2}|I_{k}^{c}\right],$
where $Z$ and $X$ are drawn from model (4) independently from $\{X_{i},Y_{i},Z_{i},D_{i}\}_{i\in I_{k}^{c}}$ and by conditioning on $I_{k}^{c}$ we mean conditioning on $\{X_{i},Y_{i},Z_{i},D_{i}\}_{i\in I_{k}^{c}}$. The analogous reasoning applies to $\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}$, $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$ and $\|l-\hat{l}^{k}\|_{L_{2}}$. The main requirements for the nuisance function estimators are that for $k=1,\ldots,K$, it holds that
$\frac{1}{\sigma_{Q}}\|l-\hat{l}^{k}\|_{L_{2}}\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right),$ (16)
$\frac{1}{\sigma_{Q}}\|l-\hat{l}^{k}\|_{L_{2}}\|f-\hat{f}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right),$
$\frac{1}{\sigma_{Q}}\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right),$
$\frac{1}{\sigma_{Q}}\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}\|f-\hat{f}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right).$
If $\sigma_{Q}^{2}$ is bounded from below, equation (16) is satisfied if all the nuisance function estimators converge at a rate of $o_{P}(N^{-1/4})$ which is comparably slow and satisfied by many modern machine learning methods, when assuming appropriate sparsity and/or smoothness conditions. If on the other hand, $\sigma_{Q}^{2}$ is decreasing with $N$, which corresponds to the setting of weak IV, the machine learners need to converge at faster rates, see also Remark 3 below.
One can show the following results (we will provide a detailed treatment for the heterogeneous treatment effect below and the results here can be obtained analogously).
Asymptotic normality of $\hat{\beta}$:
Under suitable moment conditions on the involved random variables and the functions $f$, $\phi$ and $l$ and assumptions on the nuisance function estimates including (16), we have that
$\frac{\sqrt{N}(\hat{\beta}-\beta)}{\sigma}\to\mathcal{N}(0,1)$ (17)
in distribution with
$\sigma^{2}=\frac{\mathbb{E}\left[\epsilon_{i}^{2}(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}\right]}{\mathbb{E}\left[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])\right]^{2}}.$ (18)
Consistency of $\hat{\sigma}^{2}$:
Under some additional conditions, we have that $|\hat{\sigma}^{2}/\sigma^{2}-1|=o_{P}(1)$.
Validity of robust confidence sets:
Under similar conditions, including (16), we have that under under $H_{0}:\beta=\beta_{0}$,
$\frac{\sqrt{N}\hat{Q}(\beta_{0})}{\widehat{SE}_{Q}(\beta_{0})}\to\mathcal{N}(0,1)$
in distribution, i.e. the robust procedure produces asymptotically valid confidence sets.
Remark 3.
If the nuisance functions were known, the robust confidence set construction from Section 2.3, would yield asymptotically valid inference regardless of the strength of the instruments. However, this nice property does not carry over if the nuisance functions are estimated, since in (16), the convergence rates are also scaled by the strength of the IV. The fastest convergence rate for the nuisance function estimates that we can hope for is an $L_{2}$ error of $O_{P}(N^{-1/2})$ (linear regression). Hence, in the best possible case, it still follows from (16) that $\sigma_{Q}^{2}\gg 1/N$. If we additionally assume that $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]$ is bounded away from 0 and $\infty$ independently of $N$, it follows that $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}]\asymp\sigma_{Q}^{2}\gg 1/N$. This rules out the classical “weak IV asymptotics” regime [41] which would correspond to $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}]\asymp 1/N$. The actual amount of weakness (in terms of the order of $\sigma_{Q}^{2}$) that one can in theory allow for, depends on what rates the machine learning algorithm can achieve. For example, for sparse high-dimensional linear models, one can obtain an $L_{2}$ error of $O_{P}(\sqrt{\log p/N})$ [7], so that the requirement on the IV strength is $\sigma_{Q}^{2}\gg(\log p)^{2}/N$, which is only slightly weaker than classical weak IV asymptotics. In contrast, with more flexible models such as (nonlinear) sparse high-dimensional additive models [57], neural networks [14] or (variants of) random forest [50], one can afford less weakness. In conclusion, from the analysis based on theoretical bounds one does not gain very much from using the robust confidence set construction compared to the standard confidence interval, which is in contrast to the classical treatment of weak IV. The reason for this is that compared to the classical theory, we need to estimate nuisance functions. However, for finite sample sizes and in settings with weak instruments, the robust confidence sets exhibit superior empirical coverage compared to the standard confidence intervals in simulations (see Appendix E.1).
2.5 Efficiency of the estimator $\hat{\beta}$
As detailed out in the introduction, the estimator (10) coincides with the semiparametrically efficient estimator in the homoscedastic case (see for example equation (9) in [47]). For illustration, we directly compare the asymptotic variance of $\hat{\beta}$ with the asymptotic variance of an estimator based on the identification (6). Such an estimator $\beta_{\zeta}$ is obtained by replacing $R_{f,i}^{k}$ in (10) by $R_{\zeta,i}^{k}$, where $R_{\zeta,i}^{k}=\hat{\zeta}^{k}(Z_{i},X_{i})-\hat{\mu}^{k}(X_{i})$, with $\hat{\mu}^{k}(X_{i})$ being an estimate of $\mu(X_{i})=\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$ and $\hat{\zeta}^{k}$ is either estimated or known a priori, as for example in the standard DML with linear instruments [16, 21] which is recovered by setting $\zeta(Z_{i},X_{i})=Z_{i}$. Analogously to $\hat{\beta}$, one can show that $\hat{\beta}_{\zeta}$ is asymptotically normal with asymptotic variance given by
$\sigma^{2}_{\zeta}=\frac{\mathbb{E}\left[\epsilon_{i}^{2}(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])^{2}\right]}{\mathbb{E}\left[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])\right]^{2}}.$ (19)
In the homoscedastic setting $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const.}$, the asymptotic variance of our estimator $\hat{\beta}$ is always smaller or equal the asymptotic variance of $\hat{\beta}_{\zeta}$ with equality if and only if the function $f(Z_{i},X_{i})$ is equal to the function $\zeta(Z_{i},X_{i})$ plus a potential shift $\psi(X_{i})$ depending only on $X_{i}$, as the following proposition shows.
Proposition 4.
Assume that $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathbb{E}[\epsilon_{i}^{2}]$ a.s. Then, $\sigma^{2}\leq\sigma_{\zeta}^{2}$ with equality if and only if there exist $\alpha\neq 0$ and a function $\psi$ such that
$f(Z_{i},X_{i})=\alpha\zeta(Z_{i},X_{i})+\psi(X_{i})\text{ a.s.}$ (20)
Here, $\sigma^{2}$ is as in (18), namely the asymptotic variance of the proposed estimator $\hat{\beta}$ in (10).
A proof of Proposition 4 can be found in Appendix B. The Proposition states that in the homoscedastic setting, we are estimating the optimal instrument. In particular, the estimator $\hat{\beta}$ is asymptotically more efficient than the standard estimator for double machine learning with instrumental variables [16, 21] with equal asymptotic variance if and only if there exist $\alpha\neq 0$ and a function $\psi$ such that $f(Z_{i},X_{i})=\alpha Z_{i}+\psi(X_{i})$, that is, the instrument has a linear effect on the treatment.
3 Heterogeneous Treatment Effect
Let us now consider the heterogeneous treatment effect model (1), which we restate for convenience: we assume that we observe $N$ i.i.d. realizations from
$Y_{i}=\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i}.$ (21)
The only change compared to the model (4) is that the treatment effect is assumed to be heterogeneous with respect to the univariate continuous random variable $V_{i}$ and we again assume the availability of an instrumental variable $Z_{i}\in\mathbb{R}^{d}$. As for the homogeneous treatment effect, we allow the data generating process to vary with $N$, even if it is not reflected in the notation.
The goal is to perform inference for $\beta(v)$, that is, the value of the treatment effect for a particular value of $V_{i}=v\in\mathbb{R}$. We need the following extension of Assumption 1 to make $\beta(v)$ identifiable.
Assumption 2.
  1. 1.
    $\mathbb{E}\left[\left(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}]\right)^{2}|V_{i}=v\right]\neq 0$,
  2. 2.
    $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$ ,
  3. 3.
    $V_{i}$ is measurable with respect to the sigma-algebra generated by $X_{i}$,
  4. 4.
    The distribution of $V_{i}$ is absolutely continuous with respect to the Lebesgue measure and has density $p_{V}(\cdot)$.
Assertion 1 encodes that $Z_{i}$ is a relevant IV given $X_{i}$ conditional on $V_{i}=v$. Assertion 2 is the same as for Assumption 1 and states that $Z_{i}$ is a valid instrumental variable. Assertion 3 states that $V_{i}$ is measurable with respect to $X_{i}$. Usually, $V_{i}$ is just a coordinate of $X_{i}$ and in fact, assertion 3 of Assumption 2 can be made without loss of generality since we can otherwise just replace $X_{i}\leftarrow(X_{i},V_{i})$. The absolute continuity of the distribution of $V_{i}$ with respect to the Lebesgue measure in assertion 4 is assumed for the sake of conciseness of the exposition. An analogous approach would work for discrete random variables $V_{i}$, which would reduce to simple binning. Under Assumption 2, we can identify $\beta(v)$ by simply making the identification from the homogeneous case conditional on $V_{i}=v$, i.e.
$\beta(v)=\frac{\mathbb{E}\left[(Y_{i}-\mathbb{E}[Y_{i}|X_{i}])(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])|V_{i}=v\right]}{\mathbb{E}\left[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])|V_{i}=v\right]}.$ (22)
3.1 Estimators
To construct an estimator from this identification, we use ideas from univariate kernel smoothing. For this, we need a kernel $K:\mathbb{R}\to[0,\infty)$ that satisfies the following assumption (see e.g. [51]).
Assumption 3.
  1. 1.
    $\int_{-\infty}^{\infty}K(x)\mathrm{d}x=1$,
  2. 2.
    K is symmetric, i.e. for all $x\in\mathbb{R}$, K(x)=K(x)K(x)=K(-x),
  3. 3.
    $\|K\|_{\infty}<\infty$, i.e. $K$ is bounded,
  4. 4.
    $\int_{-\infty}^{\infty}x^{2}K(x)\mathrm{d}x\in(0,\infty)$.
Consider a bandwidth $h>0$ and recall the cross-fitted residuals (7), (8) and (9). We can then modify the estimator (10) by simply weighting the products of the residuals according to the kernel, i.e.
$\hat{\beta}(v)=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)}{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)}.$ (23)
The main intuition for (23) is that we can view the treatment effect as being homogeneous when we restrict the attention to $V_{i}$ being near the target level $v$. In a similar fashion to the homogeneous treatment effect, we will show below in Section 3.3 that $\sqrt{Nh}(\hat{\beta}(v)-\hat{\beta}(v))/\sigma(v)\to\mathcal{N}(0,1)$ with asymptotic variance $\sigma^{2}(v)$ that can be consistently estimated by
$\hat{\sigma}^{2}(v)=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\hat{\beta}(v)R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}}{\left(\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}}.$ (24)
Remark 5.
As for the homogeneous treatment effect, in principle one can replace $R_{f,i}^{k}$ by $R_{\zeta,i}^{k}=\hat{\zeta}^{k}(Z_{i},X_{i})-\hat{\mu}^{k}(X_{i})$ for an arbitrary function $\zeta(Z_{i},X_{i})$ and $\hat{\mu}^{k}(X_{i})$ being an estimate of $\mu(X_{i})=\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$. Most prominently, one could take the “linear IV” estimator $\zeta(Z_{i},X_{i})=Z_{i}$ and $\mu(X_{i})=\mathbb{E}[Z_{i}|X_{i}]$. However, when using the same kernel and bandwidth, the choice $\zeta(Z_{i},X_{i})=f(Z_{i},X_{i})$ leads to the most efficient estimator of this form under homoscedasticity, as we will see in Remark 10.
3.2 Robust Confidence Sets
Also for the heterogeneous treatment effect, we can apply the robust confidence set construction. For $\gamma\in\mathbb{R}$, let
$\hat{Q}(\gamma,v)=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\gamma R_{D,i}^{k}\right)R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right),$
and
$\widehat{SE}^{2}_{Q}(\gamma,v)=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\gamma R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}-h\hat{Q}(\gamma,v)^{2}.$
Below in Section 3.3.1, we will give conditions such that under $H_{0}:\beta(v)=\beta_{0}(v)$,
$\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\widehat{SE}_{Q}(\beta_{0}(v),v)}\to\mathcal{N}(0,1),$
and consequently, $\mathcal{C}_{N}^{\alpha}(v)\coloneqq\left\{\gamma:|\hat{Q}(\gamma,v)|\leq z_{1-\alpha/2}\widehat{SE}_{Q}(\beta)/\sqrt{Nh}\right\}$ is an asymptotic $(1-\alpha)$-confidence set for $\beta(v)$, similarly to the homogeneous case.
Remark 6.
In Appendix C.1, we derive an explicit expression for the robust confidence sets and in Appendix C.2, we show that under certain conditions on the model, the robust confidence sets are asymptotically equivalent to standard confidence intervals based on $\hat{\sigma}^{2}(v)$. This means that one does not lose asymptotic efficiency by using the robust confidence sets.
3.3 Theory
In classical univariate kernel smoothing, the optimal bandwidth satisfies $h\asymp N^{-1/5}$. This yields the smallest mean squared error of the nonparametric estimator, since with this choice, the bias and the variance are of the same order. However, to obtain valid inference for $\beta(v)$, we need that the variance dominates the bias and hence, one needs to choose a smaller bandwidth (“undersmoothing”) [51]. This is also true for our setting and we need the following assumption.
Assumption 4.
$h\ll N^{-1/5}$.
The remaining assumptions can be divided into conditions on the convergence rates of the nuisance function estimates and some more technical (mainly smoothness) assumptions needed to apply the kernel smoothing successfully. We start with the assumptions on the nuisance function estimates. In contrast to the usual conditions for double/debiased machine learning that essentially require that all the nuisance functions converge faster than $N^{-1/4}$ in $L_{2}$-norm (see condition (16) above), we need slightly faster convergence requirements on the nuisance functions, which we outline below. The reason for this is that we only want to make assumptions on the overall convergence rate of the nuisance functions, e.g., $\|\hat{f}-f\|_{L_{2}}$ and not make any assumptions on convergence rates that are somehow localized/conditional on $V_{i}=v$.
Let us introduce the quantity
$\sigma_{Q}^{2}(v)=\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]p_{V}(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x,$ (25)
where we recall that $p_{V}(\cdot)$ is the density of $V_{i}$ with respect to the Lebesgue measure. Similarly to the homogeneous treatment effect, $\sigma_{Q}^{2}(v)$ can be seen as an unnormalized measure of IV strength conditional on $V_{i}=v$ on the population level.
Assumption 5.
  1. 1.
    For $k=1,\ldots,K$, it holds that $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(N^{-1/2}h^{-1}\right)$ and $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(N^{-1/2}h^{-1}\right)$.
  2. 2.
    For $k=1,\ldots,K$, it holds that $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(h^{1/2}\right)$ and $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(h^{1/2}\right)$.
  3. 3.
    For $k=1,\ldots,K$, it holds that
    $\frac{1}{\sigma_{Q}(v)}\|f-\hat{f}^{k}\|_{L_{2}}\|l-\hat{l}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(v)}\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}\|l-\hat{l}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(v)}\|f-\hat{f}^{k}\|_{L_{2}}\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(v)}\|\phi-\hat{\phi}^{k}_{1}\|_{L_{2}}\|\phi-\hat{\phi}^{k}_{2}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right).$
  4. 4.
    There exists $C<\infty$ independent of $N$ such that $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\leq C$ a.s., i.e. $\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{\infty}\leq C<\infty$.
  5. 5.
    There exist $p_{1},q_{1}\in[1,\infty]$ with $1/p_{1}+1/q_{1}=1$ and $C<\infty$ independent of $N$ such that $\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p_{1}}}/\sigma_{Q}^{2}(v)\leq C$ and for $k=1,\ldots,K$ it holds that $\|l-\hat{l}^{k}\|_{L_{2q_{1}}}=o_{P}\left(h^{1/2}\right)$.
  6. 6.
    There exist $p_{2},q_{2}\in[1,\infty]$ with $1/p_{2}+1/q_{2}=1$ and $C<\infty$ independent of $N$ such that $\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p_{2}}}/\sigma_{Q}^{2}(v)\leq C$ and for $k=1,\ldots,K$, it holds that $\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2q_{2}}}=o_{P}\left(h^{1/2}\right)$.
Remark 7.
Assume that $\sigma_{Q}^{2}(v)\asymp 1$, which is for example the case if the data generating process (21) does not change with $N$. From Assumption 4, we know that $h\ll N^{-1/5}$. Hence, assertion 1 of Assumption 5 is satisfied if $\|f-\hat{f}^{k}\|_{L_{2}}=O_{P}(N^{-3/10})$ and $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}=O_{P}(N^{-3/10})$, which is a slightly stronger requirement than $o_{P}\left(N^{-1/4}\right)$. The second assertion is comparably weak. The third assertion are the conditions on the products of the $L_{2}$ errors of the nuisance function estimators. Since $h\ll N^{-1/5}$, these assumptions are stronger than assuming that the respective products are $o_{P}\left(N^{-3/5}\right)$. Hence, the assumptions on the cross terms are also slightly stronger than for the homogeneous treatment effect.
When $\sigma_{Q}^{2}(v)\to 0$, this corresponds to models with weak IV. As already discussed in Remark 3 we cannot allow $\sigma_{Q}^{2}(v)$ to go to zero too fast since otherwise, the requirements on the convergence rates for the nuisance functions cannot be satisfied.
The remaining assumptions are more technical and we defer them to Assumption 6 in Appendix A. Assumption 6 can be summarized as follows.
  1. 1.
    $p_{V}(v)$ is bounded away from 0.
  2. 2.
    If $\sigma_{Q}^{2}(v)\to 0$ as $N\to\infty$, this is attributed to weak IV, i.e. $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]\to 0$ and not to other changes in the data generating process (for example not because $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\to 0$).
  3. 3.
    Various functions $x\mapsto\rho(x)$ are twice differentiable for $x$ in the support of $V_{i}$ with bounded second derivative (when properly rescaled by $\sigma_{Q}(v)$). The functions $\rho(\cdot)$ are products of $\beta(\cdot)$, $p_{V}(\cdot)$ and various conditional moments as for example $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=\cdot]$.
With these assumptions, we obtain the asymptotic normality of the estimator $\hat{\beta}(v)$.
Theorem 8.
Under Assumptions 2, 3, 4, 5 and 6, additionally assume that $\sigma_{Q}^{2}(v)\gg(Nh)^{-1}$, $\frac{\|f-\hat{f}^{k}\|_{L_{2}}}{\sigma_{Q}^{2}(v)}=o_{P}\left(h^{1/2}\right)$ and $\frac{\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}}{\sigma_{Q}^{2}(v)}=o_{P}\left(h^{1/2}\right)$. Then, we have
$\frac{\sqrt{Nh}(\hat{\beta}(v)-\beta(v))}{\sigma(v)}\to\mathcal{N}(0,1)\text{ in distribution,}$
where
$\sigma^{2}(v)=\frac{\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x}{\mathbb{E}[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))|V_{i}=v]^{2}p_{V}(v)}.$
The following theorem ensures that the estimator $\hat{\sigma}^{2}(v)$ defined in (24) is a consistent estimator of $\sigma^{2}(v)$ hence one can construct asymptotic confidence intervals for $\beta(v)$.
Theorem 9.
Additionally to the conditions of Theorem 8, assume that $\frac{\|f-\hat{f}^{k}\|_{L_{4}}}{\sigma_{Q}(v)}=O_{p}\left(h^{1/4}\right)$ and $\frac{\|\phi-\hat{\phi}_{2}^{k}\|_{L_{4}}}{\sigma_{Q}(v)}=O_{p}\left(h^{1/4}\right)$ Then, we have
$\frac{\hat{\sigma}^{2}(v)}{\sigma^{2}(v)}\to 1\text{ in probability.}$
Remark 10.
Regarding the efficiency of the estimator $\hat{\beta}(v)$, we can exactly follow the argument from the proof of Proposition 4. Concretely, given a fixed kernel $K$, fixed bandwidth $h$, there is no transformation $\zeta(Z_{i},X_{i})$ such that the resulting estimator according to Remark 5 is more efficient than our estimator $\hat{\beta}(v)$ under homoscedasticity $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const.}$
3.3.1 Robust Confidence Sets
The following theorem ensures that under the null hypothesis, $\hat{Q}(\beta(v),v)$ is asymptotically normal.
Theorem 11.
Under Assumptions 2, 3, 4, 5 and 6, we have that under $H_{0}:\beta(v)=\beta_{0}(v)$
$\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\sigma_{Q}(v)}\to\mathcal{N}(0,1)\text{ in distribution}.$
Moreover, the asymptotic variance of $\hat{Q}(\beta(v),v)$ can be consistently estimated, and hence, the approach for robust confidence sets is valid.
Theorem 12.
Under the assumptions of Theorem 11 we have that under $H_{0}:\beta(v)=\beta_{0}(v)$
$\frac{\widehat{SE}_{Q}^{2}(\beta_{0}(v),v)}{\sigma_{Q}^{2}(v)}\to 1\text{ in probability.}$
Remark 13.
For the validity of the robust confidence sets, one still needs Assumption 5 which does not allow for IV that are too weak (see also Remark 3). However, we do not need the additional conditions $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}^{2}(v),\,\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}/\sigma_{Q}^{2}(v)=o_{P}(h^{1/2})$ that – given a rate of convergence for $\hat{f}^{k}$ and $\hat{\phi}_{2}^{k}$ – pose even stronger restrictions on the IV-strength $\sigma_{Q}^{2}(v)$. In that sense, the robust confidence set construction works under slightly less restrictive conditions than the asymptotic normality of the point estimator.
4 Applications
We evaluate our method on simulated and real-world datasets. The method is implemented in the R package IVDML available on CRAN [38]. The code to reproduce the simulation and real data results is available on GitHub. [^1] https://github.com/cyrillsch/IVDML_Application Before considering simulations and real data applications, we point out an aspect that is important when applying the methods in practice.
4.1 Repeated Cross-Fitting
The partition of the dataset into the $K$ folds is random. The finite sample estimators $\hat{\beta}$ and $\hat{\sigma}^{2}$ for the homogeneous treatment effect and $\hat{\beta}(v)$ and $\hat{\sigma}^{2}(v)$ for the heterogeneous treatment effect depend on this random partition. Let us focus on the homogeneous treatment effect in the following (the aggregation for the heterogeneous treatment effect is analogous). Following [16], the procedure can be repeated $S$ times with independent random partitions, yielding estimators $(\beta_{s})_{s=1,\ldots,S}$ and $(\hat{\sigma}^{2}_{s})_{s=1,\ldots,S}$. The point estimator is then aggregated using the median, $\hat{\beta}^{*}=\mathrm{median}\{\hat{\beta}_{s};\,s=1,\ldots,S\}$ and the asymptotic variance additionally is updated by a term incorporating the sample variability, $\hat{\sigma}^{2,*}=\mathrm{median}\{\hat{\sigma}_{s}^{2}+(\hat{\theta}_{s}-\hat{\theta}^{*})^{2};\,s=1,\ldots,S\}$. For the robust confidence set construction, we use the same idea. Let $\left(\hat{Q}_{s}(\cdot)\right)_{s=1,\ldots,S}$ and $\left(\widehat{SE}_{Q,s}^{2}(\cdot)\right)_{b=1,\ldots,B}$ be obtained from (13) and (14). Define $\hat{Q}^{*}(\cdot)=\mathrm{median}\left\{\hat{Q}_{s}(\cdot);\,s=1,\ldots,S\right\}$ and $\widehat{SE}_{Q}^{2,s}(\cdot)=\mathrm{median}\left\{\widehat{SE}_{Q,s}^{2}(\cdot)+\left(\hat{Q}_{s}(\cdot)-\hat{Q}^{*}(\cdot)\right)^{2};\,s=1,\ldots,S\right\}$, where the medians are taken argument-wise.
4.2 Simulations
We simulate from model (21) using the following specification:
$X_{i},H_{i},E_{Z,i},E_{\delta,i},E_{\epsilon,i}$ $\sim_{i.i.d.}\mathcal{N}(0,1),$ (26)
[…]
$D_{i}$ $\leftarrow f(Z_{i},X_{i})+\delta_{i},$
$V_{i}$ $\leftarrow X_{i},$
$Y_{i}$ $\leftarrow\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i}.$
In (26), $H_{i}$ acts as unobserved confounding, which is the reason we need the instrumental variable $Z_{i}$. We fix $g(x)=\tanh(x)$ and consider combinations of the following settings for $f$ and $\beta$:
$\beta(v)=1$ (homogeneous treatment effect),
$\beta(v)=2\exp(-v^{2}/2)$ (heterogeneous treatment effect),
(Z lin.)
$f(z,x)=-\sin(x)+z$ (instrument acts linearly on treatment),
(Z nonlin.)
$f(z,x)=-\sin(x)+\cos(z)+0.2z$ (instrument acts nonlinearly on treatment).
Moreover, we fit the following models, where we use 5-fold cross-fitting repeated $S=10$ times and aggregate the results using the method described in Section 4.1.
(hom. T.E. mlIV)
The DML estimator, see (10) and Remark 2 for the homogeneous treatment effect with machine learning instruments.
(hom. T.E. linearIV)
The original DML estimator for the homogeneous treatment effect based on (5) using the instrument $Z_{i}$ linearly [16, 21].
(het. T.E. mlIV)
The estimator (23) for the heterogeneous treatment effect with machine learning instruments. Depends on the choice of a bandwidth and a kernel.
(het. T.E. linearIV)
Variant of the estimator (23) for the heterogeneous treatment that uses the instruments only linearly. Depends on the choice of a bandwidth and a kernel.
In all the following simulations, we estimate the nuisance functions using generalized additive models with the function gam from the R package mgcv [52, 53] using the default settings. As kernel, we use the Epanechnikov kernel [51], $K(x)=\frac{3}{4\sqrt{5}}(1-x^{2}/5)1\{|x|\leq\sqrt{5}\}$, which is scaled such that $\int x^{2}K(x)\mathrm{d}x=1$.
First, we simulate two datasets with $N=1000$ i.i.d. samples, one from setting (het.)/(Z lin.) and one from setting (het.)/(Z nonlin.) and apply the methods (het. T.E. mlIV) and (het. T.E. linearIV). We use two different bandwidths. The first one is the normal reference rule / Silverman’s rule of thumb [40, 51] and is given by $h_{N}=1.06\min(s,Q/1.34)N^{-1/5}$, where $s$ is the sample standard deviation and $Q$ the interquartile range of $\{V_{i}\}_{i=1}^{N}$. For valid inference, we need undersmoothing $h_{N}\ll N^{-1/5}$ (Assumption 4) and hence for the smaller bandwidth, we replace the $N^{-1/5}$ by $N^{-2/7}$ in the normal reference rule, which is also, what is done in [24]. The estimated heterogeneous treatment effect functions together with (pointwise) standard confidence intervals (dashed lines) and (pointwise) robust confidence sets (shaded region) can be seen in Figure 1. From the plots, we see that in the setting (Z lin.), the methods (het. T.E. mlIV) and (het. T.E. linearIV) yield almost identical results, both in terms of point estimate and confidence sets. In contrast, in the setting (Z nonlin.), the method (het. T.E. mlIV) seems to yield more stable results with smaller confidence bounds, whereas the results using (het. T.E. linearIV) are very unstable even in regions where there are a lot of samples $V_{i}$. This is consistent with our theory (see Remarks 5 and 10). Moreover, we also observe that the standard confidence intervals and the robust confidence sets are almost identical in regions with a lot of samples $V_{i}$ and large instrument strength. In contrast, if either there are not a lot of samples $V_{i}$ near a point $v$ (e.g. at the boundary) or the strength of the instrument is not very strong (e.g. method (het. T.E. linearIV) in the setting (Z nonlin.)), the robust confidence sets can be considerably larger than the standard confidence sets.
Figure 1: Visualization for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom). On the left, the bandwidth $h$ is chosen according to the normal reference rule / Silverman’s rule of thumb. On the right, the bandwidth is chosen according to the normal reference rule multiplied by $N^{1/5}/N^{2/7}$ to ensure undersmoothing. The black solid line is the true heterogeneous treatment effect. Solid lines are the point estimates for the methods (het. T.E. linearIV) (red) and (het. T.E. mlIV) (blue). Dashed lines are the (pointwise) standard 95%-confidence interval. Shaded regions are the (pointwise) robust 95%-confidence sets. Red-shaded regions correspond to (het. T.E. linearIV) and blue-shaded regions correspond to (het. T.E. mlIV). Purple-shaded regions are the intersection of the two.
We now want to investigate the efficiency of the different estimators and the coverage of the different confidence sets more extensively. For each combination of setting (hom.)/(het.) with (Z lin.)/(Z nonlin.) we consider sample size $N\in\{250,500,1000,1500,2000\}$ and simulate $1000$ datasets from (26). We still use generalized additive models (gam) for the nuisance functions and the Epanechnikov kernel for methods (het. T.E. mlIV) and (het. T.E. linearIV). Moreover, we only consider undersmoothing for the bandwidth, i.e. we use the normal reference rule times $N^{1/5}/N^{2/7}$. Let us first consider the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.). Here, we use the methods (hom. T.E. mlIV) and (hom. T.E. linearIV) as well as the methods (het. T.E. mlIV) and (het. T.E. linearIV) for $\hat{\beta}(0)$ and $\hat{\beta}(1.5)$. The results are shown in Figure 2. As it is to be expected, the mean squared errors (MSEs) of the estimators decrease with $N$ and the MSEs of the (hom. T.E.) estimators are smaller than the MSEs of the comparable (het. T.E.) estimators. The MSEs of the (het. T.E.) estimators for $v=0$ are smaller than the MSEs for the (het. T.E.) estimators for $v=1.5$, which is due to the fact that the density of $V_{i}$ is larger at $v=0$ than at $v=1.5$. Moreover, we see that both the standard and robust confidence sets exhibit good empirical coverage close to the desired 0.95. Regarding the difference between settings (Z lin.) and (Z nonlin.), we see that when the instrument $Z$ acts linearly on the treatment (Z lin.), the performance of linearIV and mlIV is very similar for both the homogeneous treatment effect and the heterogeneous treatment effect. On the other hand, we observe that in the setting (Z nonlin.), the methods based on mlIV have significantly lower MSE than their corresponding linearIV counterpart. This reflects the theoretical results that using the optimal machine learning instruments can lead to a significant efficiency gain.
Figure 2: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom). Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Let us now consider the settings (het.)/(Z lin.) and (het.)/(Z nonlin.). Here, we only fit the methods that actually estimate a heterogeneous treatment effect. The results are shown in Figure 3. The overall picture is similar to before. However, the coverage of the standard confidence interval for the (het. T.E.) methods at $v=1.5$ are a bit too low, but the coverage of the robust confidence sets is better. In the setting (Z nonlin.), the methods based on mlIV have again lower MSEs than the methods based on linearIV.
Figure 3: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom). Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
In Appendix E, we present additional simulations: in Section E.1, we vary the IV strength and see that the standard confidence intervals may suffer from substantial undercoverage, whereas the robust confidence intervals still exhibit coverage close to the desired level of 0.95. In Section E.2, we consider higher-dimensional covariates and nuisance function estimation using boosted regression trees instead of generalized additive models.
4.3 Real Data Experiments
We apply our methods to two well-known instrumental variables datasets. Additionally to gam (generalized additive models using the R package mgcv [52, 53]), we also use boosted regression trees and random forest to estimate the nuisance functions. We use the R packages xgboost [12, 13] and ranger [55]. We refer to the three machine learning methods as gam, xgboost and random forest. In contrast to gam, for which we simply use the default settings, the methods xgboost and random forest need to be tuned. In view of the cross-fitting scheme, one would need to tune the machine learning methods for each of the $K$ different training datasets of size NN/KN-N/K separately to avoid cross-contamination of the sample. However, recent research by [5] showed that this is usually not an issue and it suffices to tune the machine learning methods once on the full sample. Hence, in the following, for each nuisance function that has to be estimated, we tune the machine learner once on the full sample to obtain the optimal hyperparameters (with respect to cross-validated prediction accuracy for xgboost and with respect to out-of-bag prediction error for random forest) and then keep them fixed for the $K$ subsamples in all $S$ repetitions of the aggregation procedure from Section 4.1.
4.3.1 Effect of Institutions on Economic Growth
We revisit the question of the effect of institutions on economic performance. The dataset originates from [1] and is available from the R-package hdm [15]. In the following, we refer to it as the AJR dataset. It was previously analyzed in [16] and [21]. The AJR dataset contains 64 country-level observations of the GDP, an index measuring protection against expropriation, settler mortality, and geographic information. In our notation, the response $Y$ is the logarithm of GDP per capita, $D$ is the expropriation protection index, the instrument $Z$ is the logarithm of settler mortality and $X$ contains the latitude and dummies for the continents (Africa, Asia, North America and South America). While the relationship between expropriation protection and GDP might be endogenous, it is argued that settler mortality is a valid instrument, since it is unlikely that settler mortality has an influence on today’s GDP other than through institutions, especially when controlling for geographic information. We refer to [1, 16] for a more detailed discussion of the dataset.
We first apply the methods (hom. T.E. linearIV) and (hom. T.E. mlIV) to the dataset. We use 5-fold cross-fitting with $S=200$ repetitions in the procedure from Section 4.1 and for the nuisance function estimates, we use gam, xgboost and random forest as described in the previous subsection. The results can be found in Table 1. Point estimates, standard deviations and standard confidence intervals were also considered in [16, 21] and our results are consistent with their findings up to variability due to sampling and tuning of the machine learning methods (the sample size is only $N=64$). The robust confidence set based on (linearIV) was also considered in Section 13.3 of [17]. What is new here, are the results based on (mlIV). We see that in most of the cases, the use of (mlIV) compared to (linearIV) leads to a reduction of the standard deviation of the estimator, the length of the standard confidence interval and the size of the robust confidence sets. We also observe that the robust confidence sets can be disconnected. Recall from Section 2.3 that the robust confidence set is defined as the sublevel set of a parabola, which can be disconnected if the parabola has positive curvature.
$\hat{\beta}$ $\hat{\sigma}/\sqrt{n}$ standard CI robust CI
linearIV (gam) 0.72 0.27 $[0.20,1.25]$ $[0.29,4.02]$
mlIV (gam) 0.58 0.16 $[0.27,0.90]$ $[0.28,1.81]$
linearIV (xgboost) 0.62 0.35 [0.07,1.31][-0.07,1.31] $(-\infty,\infty)$
mlIV (xgboost) 1.05 0.35 $[0.36,1.75]$ $(-\infty,-3.84]\cup[0.47,\infty)$
linearIV (random forest) 0.84 0.43 $[0.00,1.68]$ $(-\infty,-0.50]\cup[0.04,\infty)$
MLIV (random forest) 0.94 0.36 $[0.24,1.65]$ $(-\infty,-3.90]\cup[0.25,\infty)$
Table 1: Results for the methods (hom. T.E. linearIV) and (hom. T.E. mlIV) on the AJR dataset.
One might wonder if the homogeneous treatment effect is a realistic model assumption for the AJR dataset or if one should consider heterogeneity with respect to geography. We do not aim to give a comprehensive answer to this question, but for illustration purposes, we apply the methods (het. T.E. mlIV) and (het. T.E. linearIV), where we take the latitude as the variable $V_{i}$ with respect to which we want to quantify the heterogeneity. In Figure 4, we see the estimated heterogeneous effects with respect to latitude for the machine learning methods gam and xgboost. We omit the results for random forest as they look qualitatively similar. As for the initial example in Section 4.2, we use the Epanechnikov kernel and the bandwidth according to the normal reference rule as well as a smaller bandwidth to ensure undersmoothing. From the plots, we cannot infer that there is a significant heterogeneity with respect to latitude. In fact, the robust confidence sets are unbounded for most values of latitude for both (linearIV) and (mlIV), especially when using xgboost for the nuisance functions and the smaller bandwidth. Note that due to the large difference between the robust confidence sets and the standard confidence intervals, we should not trust the standard confidence intervals. Especially, the size of the standard confidence intervals seems to decrease for large latitude. It is important to note, that this is an artifact of the data. For example, the confidence interval with latitude around $0.7$ is only based on the two datapoints with latitude close to $0.7$. Hence, there is no substantive value in the size of those confidence intervals. We conclude that for the AJR dataset, the sample size of $N=64$ is too small to make a sensible statement about a potential heterogeneous treatment effect.
Figure 4: Heterogeneous treatment effect for the AJR data. Top: nuisance functions are estimated using gam. Bottom: nuisance functions are estimated using xgboost. On the left, the bandwidth $h$ is chosen according to the normal reference rule / Silverman’s rule of thumb. On the right, the bandwidth is chosen according to the normal reference rule multiplied by $N^{1/5}/N^{2/7}$ to ensure undersmoothing. Solid lines are the point estimates for the methods (het. T.E. linearIV) (red) and (het. T.E. mlIV) (blue). Dashed lines are the (pointwise) standard 95%-confidence interval. Shaded regions are the (pointwise) robust 95%-confidence sets. For the smaller bandwidth, there is a region around 0.57, for which no results can be estimated because no datapoints are close enough. Red-shaded regions correspond to (het. T.E. linearIV) and blue-shaded regions correspond to (het. T.E. mlIV). Purple-shaded regions are the intersection of the two.
4.3.2 Effect of Education on Wage
We consider the famous dataset by [10] on the effect of education on wage. The dataset is available from the R-package ivmodel [28]. We refer to the dataset as the Card dataset. It contains data for $N=3010$ individuals from the National Longitudinal Survey of Young Men. We follow the recent analysis by [26] and use the same treatment, response, instrument and covariates as there. Concretely, in our notation, the response $Y_{i}$ is the logarithm of the wage of individual $i$, $D_{i}$ is the years of education of individual $i$, the instrument $Z_{i}$ is an indicator of whether the individual grew up near a four-year college and $X_{i}$ are covariates, including years of experience and binary indicators for race and geographic information. The idea for using college proximity as an instrument for education is that, while there are factors that influence both education and wage, they should be independent of whether an individual grew up near a college or not. As before, we first apply the methods (hom. T.E. linearIV) and (hom. T.E. mlIV) to the dataset. We use 5-fold cross-fitting with $S=50$ repetitions for the procedure from Section 4.1. For the nuisance function estimates, we use gam, xgboost and random forest. The results can be found in Table 2. In contrast to the AJR dataset, we see that the use of (mlIV) compared to (linearIV) does not lead to a reduction of the standard deviation of the estimator, the length of the standard confidence interval and the size of the robust confidence sets. It seems that on this dataset, there is nothing to gain from using (mlIV) and one loses a bit of performance (i.e. larger confidence intervals) due to the additional variability that is introduced in the procedure by having to learn the instruments instead of using them linearly.
$\hat{\beta}$ $\hat{\sigma}/\sqrt{n}$ standard CI robust CI
linearIV (gam) 0.14 0.05 $[0.04,0.24]$ $[0.05,0.28]$
mlIV (gam) 0.14 0.05 $[0.03,0.24]$ $[0.03,0.28]$
linearIV (xgboost) 0.12 0.04 $[0.04,0.21]$ $[0.04,0.23]$
mlIV (xgboost) 0.11 0.05 $[0.02,0.20]$ $[0.01,0.23]$
linearIV (random forest) 0.13 0.05 $[0.04,0.23]$ $[0.05,0.25]$
MLIV (random forest) 0.14 0.08 [0.01,0.3][-0.01,0.3] $[0.01,0.60]$
Table 2: Results for the methods (hom. T.E. linearIV) and (hom. T.E. mlIV) on the Card dataset.
Also for the Card dataset, we apply the methods (het. T.E. mlIV) and (het. T.E. linearIV) to see if we can detect heterogeneity of the treatment effect. As the variable $V_{i}$, we take the years of experience of individual $i$. The estimated heterogeneous treatment effects with respect to experience are displayed in Figure 5. Again, we only show the results for the machine learning methods gam and xgboost and omit the results for random forest since they look similar. Moreover, we still use the Epanechnikov kernel and the bandwidth according to the normal reference rule as well as a smaller bandwidth to ensure undersmoothing.
Similarly to the AJR datasets, we cannot infer a significant heterogeneity from the plots. Moreover, we see that the methods based on (mlIV) have larger confidence intervals, in particular when the nuisance functions are estimated using xgboost. Interestingly, there are some regions for experience, where the value of the heterogeneous treatment effect seems to be quite narrowly determined. On the other hand, there are regions, where there is a lot of uncertainty in the value of the heterogeneous treatment effect (indicated by larger standard and robust confidence sets). Surprisingly, those effects do not only happen in low-density regions of experience (near the boundary) but also around 10 years of experience, where the density of experience is rather high. Unfortunately, it is hard to determine, if this is an artifact of the sample or indeed has a substantive meaning.
Figure 5: Heterogeneous treatment effect for the Card data. Top: nuisance functions are estimated using gam. Bottom: nuisance functions are estimated using xgboost. On the left, the bandwidth $h$ is chosen according to the normal reference rule / Silverman’s rule of thumb. On the right, the bandwidth is chosen according to the normal reference rule multiplied by $N^{1/5}/N^{2/7}$ to ensure undersmoothing. Solid lines are the point estimates for the methods (het. T.E. linearIV) (red) and (het. T.E. mlIV) (blue). Dashed lines are the (pointwise) standard 95%-confidence interval. Shaded regions are the (pointwise) robust 95%-confidence sets. Red-shaded regions correspond to (het. T.E. linearIV) and blue-shaded regions correspond to (het. T.E. mlIV). Purple-shaded regions are the intersection of the two.
5 Discussion
We developed a novel method to perform inference on potentially heterogeneous treatment effects using double machine learning and efficient machine learning instruments. Our method is simple and yields results that are easy to interpret. We gave conditions under which our estimator is asymptotically Gaussian and also constructed confidence sets that are robust to a certain extent of weak instruments. Additionally, we hope that our paper is also helpful for readers who are looking for an accessible discussion of the DML estimator with machine learning instruments for the homogeneous treatment effect.
An important avenue for further research are methods to decide from data if a constant treatment effect model is enough or if a heterogeneous treatment effect model is more appropriate. As an informal test, one can check if the homogeneous treatment effect estimate (i.e., a horizontal line) lies inside the confidence sets for the heterogeneous treatment effect. In fact, from the corresponding plots in our two real-data case studies, we observed that this is (more or less) the case which leads to the conclusion that the heterogeneous treatment effect model is probably not needed. However, the confidence sets that we provided are only pointwise. Hence, it would also be interesting to have uniform confidence bands. We expect that this is possible using multiplier bootstrap, as for example done in [24].
[…]
A limitation of our approach is that it only considers heterogeneity with respect to a univariate quantity, i.e. the dimension of $V_{i}$ is $1$. If the treatment effect is potentially heterogeneous with respect to the full set of covariates $X_{i}$, i.e. the true model is $Y_{i}=\beta(X_{i})D_{i}+g(X_{i})+\epsilon_{i}$, equation (22) identifies $\mathbb{E}[\beta(X_{i})w(Z_{i},X_{i},v)|V_{i}=v]$ with the weight function
$w(Z_{i},X_{i},v)=\frac{(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}}{\mathbb{E}[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}|V_{i}=v]}.$
In particular, this is not just the projection $\mathbb{E}[\beta(X_{i})|V_{i}=v]$ of the true heterogeneous treatment effect on $V_{i}$.
A straightforward extension of our approach would be to use multivariate kernel smoothing to deal with multivariate $V_{i}$ and the results should carry over. However, note that the conditions on the convergence rates of the nuisance function estimators (Assumption 5) depend on the bandwidth $h$. In higher dimensions, they will depend on $h^{d}$, where $h$ is the dimension of $h$ and it will quickly become unrealistic to have machine learning methods that achieve these convergence rates.
We can also take one step back. In principle, we can replace the outer conditional expectations in (22) by arbitrary machine learning methods potentially also allowing for higher-dimensional $V_{i}$. A natural assumption might be that $V_{i}=X_{i}$ which is potentially high-dimensional, but that the treatment effect is only heterogeneous with respect to a small subset of the components of $X_{i}$, so one could use methods enforcing sparsity for the outer conditional expectation. If the machine learning method in the numerator allows for inference and the machine learning method in the denominator is consistent, one might still be able to perform inference for the heterogeneous treatment effect.
The aim of this work was to present a simple and ready-to-use method for inference in combination with user-chosen machine learning algorithms, but from a theoretical perspective, we did not strive for full generality. For the homogeneous case, estimators with better theoretical properties could probably be obtained using higher order influence functions [36] or triple sample splitting [32]. For the heterogeneous treatment effect – at least in the case of binary treatment and without endogeneity – minimax optimality was discussed by [29, 30]. Whereas this was not the focus of our work, it would be interesting to obtain similar results for the endogenous setting and allowing for continuous treatment as considered here.
We leave all these interesting directions to further research.
Acknowledgements
We are grateful to Malte Londschien for helpful discussions about weak instrumental variables. CS received funding from the Swiss National Science Foundation, grant no. 214865. ZG acknowledges financial support for visiting the Institute of Mathematical Research (FIM) at ETH Zurich. ZG was partly supported by NIH R01LM013614, NIH R01AG086379, and NSF DMS 2413107.
Appendix A Technical Conditions for the Heterogeneous Treatment Effect
The following assumptions are needed to be able to estimate and perform inference for the heterogeneous treatment effect $\beta(v)$.
Assumption 6.
  1. 1.
    The density $p_{V}(\cdot)$ of $V_{i}$ with respect to Lebesgue measure satisfies that $p_{V}(v)\geq C$ for a constant $C>0$ independent of $N$.
  2. 2.
    $\sigma_{Q}^{2}(v)\asymp\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]$.
  3. 3.
    $\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v\right]\lesssim\sigma_{Q}^{2}(v)$.
  4. 4.
    The function $\rho(x)=(\beta(x)-\beta(v))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}(v)<C$.
  5. 5.
    The function $\rho(x)=\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
  6. 6.
    The function $\rho(x)=(\beta(x)-\beta(v))\mathbb{E}[\epsilon_{i}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
  7. 7.
    The function $\rho(x)=(\beta(x)-\beta(v))^{2}\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
  8. 8.
    The function $\rho(x)=(\beta(x)-\beta(v))^{2}\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}<C$.
  9. 9.
    The function $\rho(x)=\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
  10. 10.
    The function $\rho(x)=\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
  11. 11.
    The function $\rho(x)=\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}<C$.
  12. 12.
    There exists $\eta>0$ and $C>0$ independent of $N$ such that
    $\frac{\mathbb{E}\left[\left|\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)\right|^{2+\eta}\right]}{\sigma_{Q}(v)^{2+\eta}}\leq Ch.$
  13. 13.
    $\frac{\mathbb{E}\left[\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)^{4}\left(f(Z_{i},X_{i})-\phi(X_{i})\right)^{4}K\left(\frac{V_{i}-v}{h}\right)^{4}\right]}{\sigma_{Q}^{4}(v)}\ll Nh^{2}.$
  14. 14.
    $\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{4}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]=O(h).$
Remark 14.
Note that all the quantities in Assumption 6 may depend on $N$ even if it is not reflected in the notation. Assertion 1 intuitively means that we need to have enough samples $i$ with $V_{i}$ close to $v$ to be able to estimate the heterogeneous treatment effect. Assertions 2 and 3 are needed in the case where the IV strength $\sigma_{Q}^{2}(v)$ is decreasing. Essentially they state that a decreasing $\sigma_{Q}^{2}(v)$ is attributed to $f(Z_{i},X_{i})$ being close to $\phi(X_{i})$. They are for example satisfied if $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]$ and $\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|Z_{i},X_{i}]$ are both bounded away from zero and infinity independent of $N$. Assertions 4-11 are smoothness assumptions on $\beta(x)$, $p_{V}(x)$ and conditional moments of $\epsilon_{i}$, $D_{i}$, $f(Z_{i},X_{i})$ and $\phi(X_{i})$. These smoothness conditions need to hold independently of $N$. In view of Lemma 23 in Appendix D.5 (with the kernel $K$ potentially replaced by $K^{2+\eta}$ or $K^{4}$), assertions 12-14 are comparably weak and hold in particular, if the quantities inside the expectations satisfy similar smoothness conditions as assertions 4-11.
Appendix B Proof of Proposition 4
Applying the tower property for conditional expectations to (18), we have that
$\sigma^{2}=\frac{\mathbb{E}\left[\epsilon_{i}^{2}(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}\right]}{\mathbb{E}\left[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}\right]^{2}}.$
Since by assumption, $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathbb{E}[\epsilon_{i}^{2}]$, we can simplify
$\sigma^{2}$ $=\frac{\mathbb{E}[\epsilon_{i}^{2}]}{\mathbb{E}\left[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}\right]},$
$\sigma^{2}_{\zeta}$ $=\frac{\mathbb{E}\left[\epsilon_{i}^{2}\right]\mathbb{E}\left[(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])^{2}\right]}{\mathbb{E}\left[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])\right]^{2}}.$
For the denominator of $\sigma^{2}_{\zeta}$, we have by the Cauchy-Schwarz inequality
$\mathbb{E}\left[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])\right]^{2}$
$\leq\mathbb{E}\left[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}\right]\mathbb{E}\left[(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])^{2}\right].$
It follows that $\sigma^{2}\leq\sigma^{2}_{\zeta}$ with equality if and only if there exists $\alpha\neq 0$ such that $\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}]=\alpha(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])$ a.s. Now, observe that if $\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}]=\alpha(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}])$, (20) holds with $\psi(X_{i})=\mathbb{E}[D_{i}|X_{i}]-\alpha\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$. If on the other hand (20) holds, we have
$\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}]$ $=\alpha\zeta(Z_{i},X_{i})+\psi(X_{i})-\alpha\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]-\psi(X_{i})$
$=\alpha(\zeta(Z_{i},X_{i})-\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]).$
This concludes the proof.
Appendix C More on the Robust Confidence Set
C.1 Explicit Expression for the Robust Confidence Set
Using basic manipulations, we can derive explicit expressions for $\mathcal{C}_{N}^{\alpha}(v)$. For this, define the quantities
$A_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$
$B_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$
$C_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$E_{N}$ $=\frac{2}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{D,i}^{k}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$F_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}.$
Then, we have that
$\mathcal{C}_{N}^{\alpha}(v)=\left\{\gamma:(A_{N}-\gamma B_{N})^{2}\leq\frac{z_{1-\alpha/2}^{2}}{Nh}\left(C_{N}-\gamma E_{N}+\gamma^{2}F_{N}-h(A_{N}-\gamma B_{N})^{2}\right)\right\}$
or equivalently $\mathcal{C}_{N}^{\alpha}(v)=\left\{\gamma:R_{N}\gamma^{2}+S_{N}\gamma+T_{N}\leq 0\right\}$ with
$R_{N}$ $=B_{N}^{2}+\frac{z_{1-\alpha/2}^{2}}{Nh}(hB_{N}^{2}-F_{N})$
$S_{N}$ $=-2A_{N}B_{N}+\frac{z_{1-\alpha/2}^{2}}{Nh}(E_{N}-2hA_{N}B_{N})$
$T_{N}$ $=A_{N}^{2}+\frac{z_{1-\alpha/2}^{2}}{Nh}(hA_{N}^{2}-C_{N})$
Hence, we have the following explicit expression
$\mathcal{C}_{N}^{\alpha}(v)=\begin{cases}\left[\frac{-S_{N}-\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}},\frac{-S_{N}+\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}}\right],&R_{N}>0\\ \mathbb{R}\setminus\left[\frac{-S_{N}+\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}},\frac{-S_{N}-\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}}\right],&R_{N}<0.\end{cases}$ (27)
By construction, we have that $\hat{\beta}(v)=A_{N}/B_{N}\in\mathcal{C}_{N}^{\alpha}(v)$. Hence, in the case $R_{N}>0$, it will always hold that $S_{N}^{2}-4R_{N}T_{N}>0$. In the case $R_{N}<0$, it may happen that $S_{N}^{2}-4R_{N}T_{N}<0$, in which case, $\mathcal{C}_{N}^{\alpha}(v)=\mathbb{R}$.
C.2 Adaptivity of the Robust Confidence Set
Proposition 15.
Assume that there exist $A_{0},B_{0},C_{0},E_{0},F_{0}\in\mathbb{R}$ with $B_{0}\neq 0$ such that $A_{N}=A_{0}+o_{P}(1)$, $B_{N}=B_{0}+o_{P}(1)$, $C_{N}=C_{0}+o_{P}(1)$, $E_{N}=E_{0}+o_{P}(1)$ and $F_{N}=F_{0}+o_{P}(1)$. Then, the confidence set $\mathcal{C}_{N}^{\alpha}(v)$ is asymptotically equivalent to the standard confidence interval, i.e. for fixed $\alpha\in(0,1)$,
$C_{N}^{\alpha}(v)=\left[\hat{\beta}(v)-\frac{z_{1-\alpha/2}\hat{\sigma}(v)}{\sqrt{Nh}}+o_{P}\left(\frac{1}{\sqrt{Nh}}\right),\,\hat{\beta}(v)+\frac{z_{1-\alpha/2}\hat{\sigma}(v)}{\sqrt{Nh}}+o_{P}\left(\frac{1}{\sqrt{Nh}}\right)\right]$
Remark 16.
If the data generating process is fixed, then under the assumptions of Theorem 8, one can easily show that the conditions of Proposition 15 are satisfied by using the same arguments as in the proof of Theorem 8.
Proof of Proposition 15.
We use expression (27). Since $B_{N}=B_{0}+o_{P}(1)$, $F_{N}=F_{0}+o_{P}(1)$ and $B_{0}\neq 0$, we have that $\mathbb{P}(R_{N}>0)\to 1$ and we only need to consider the case $R_{N}>0$ in (27). Hence, we can write
$\mathcal{C}_{N}^{\alpha}(v)=\left[\frac{A_{N}B_{N}\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{B_{N}^{2}C_{N}+A_{N}^{2}F_{N}-A_{N}B_{N}E_{N}+O_{P}\left(\frac{1}{Nh}\right)}+O_{P}\left(\frac{1}{Nh}\right)}{B_{N}^{2}+O_{P}\left(\frac{1}{Nh}\right)}\right].$
Because $B_{N}^{2}=B_{0}^{2}+o_{P}(1)$ and using that $|\sqrt{a+b}-\sqrt{a}|\leq\sqrt{|b|}$ for $a>0$ and $b\geq-a$ it follows that
$\textstyle\mathcal{C}_{N}^{\alpha}(v)$ $=\textstyle\left[\frac{A_{N}}{B_{N}}\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{\frac{C_{N}}{B_{N}^{2}}+\frac{A_{N}^{2}F_{N}}{B_{N}^{4}}-\frac{A_{N}E_{N}}{B_{N}^{3}}+O_{P}\left(\frac{1}{Nh}\right)}+O_{P}\left(\frac{1}{Nh}\right)\right]$
$=\textstyle\left[\hat{\beta}(v)\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{\frac{C_{N}}{B_{N}^{2}}+\hat{\beta}(v)^{2}\frac{F_{N}}{B_{N}^{2}}-\hat{\beta}(v)\frac{E_{N}}{B_{N}^{2}}}+O_{P}(1/(Nh))\right]$
where by (23), $\hat{\beta}(v)=A_{N}/B_{N}$. Note that
$\frac{C_{N}}{B_{N}^{2}}+\hat{\beta}(v)^{2}\frac{F_{N}}{B_{N}^{2}}-\hat{\beta}(v)\frac{E_{N}}{B_{N}^{2}}=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(R_{Y,i}^{k}-\hat{\beta}(v)R_{D,i}^{k})^{2}(R_{f,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}}{\left(\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}},$
which is equal to $\hat{\sigma}^{2}(v)$ defined in (24). Hence, the robust and the non-robust confidence intervals are asymptotically equivalent. ∎
Appendix D Proofs for Section 3
The proofs use similar ideas as [16] combined with standard ideas from univariate kernel smoothing.
D.1 Proof of Theorem 11
The strategy is to replace the estimated nuisance functions $\hat{f}^{k},\hat{\phi}_{1}^{k},\hat{\phi}_{2}^{k},\hat{l}^{k}$ by the true functions $f,\phi,l$ in $\hat{Q}(\beta_{0}(v),v)$. In fact, the theorem follows from the following two lemmas.
Lemma 17.
Under the conditions of Theorem 11,
$\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\to\mathcal{N}(0,1),\text{ in distribution.}$
Lemma 18.
Under the conditions of Theorem 11,
$\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle-\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\sigma_{Q}(v)}=o_{P}(1).$
D.1.1 Proof of Lemma 17
Under $H_{0}$, we have $\beta(v)=\beta_{0}(v)$ and hence $Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i}))=\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))$. Hence, we need to consider the i.i.d. random variables
$W_{i}=\frac{\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)}{\sqrt{h}\sigma_{Q}(v)}.$
By Assumption 2, $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$. We can use the tower property of conditional expectations together with the fact that $V_{i}$ is measurable with respect to $X_{i}$, $\mathbb{E}[D_{i}|Z_{i},X_{i}]=f(Z_{i},X_{i})$ and Lemma 23 from Section D.5 below to arrive at
$\mathbb{E}[W_{i}]$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[(\beta(V_{i})-\beta(v))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}]K\left(\frac{V_{i}-v}{h}\right)\right]$
$=h^{2.5}\frac{\rho^{\prime\prime}(v)}{2\sigma_{Q}(v)}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x+\frac{R(v,h)}{\sqrt{h}\sigma_{Q}(v)}.$
with $\rho(x)=(\beta(x)-\beta(v))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ and $R(v,h)$ defined accordingly in Lemma 23. By assertion 4 of Assumption 6, there exists a constant $C\in\mathbb{R}$ independent of $N$ such that
$E[W_{i}]\leq Ch^{5/2}.$ (28)
Similarly, from the second assertion of Lemma 23 together with assertions 5, 6 and 7 of Assumption 6
$\textstyle\mathbb{E}[W_{i}^{2}]=\textstyle\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$
$\quad\textstyle+\frac{2}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\epsilon_{i}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$
$\quad\textstyle+\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$
$\textstyle=\frac{1}{\sigma_{Q}^{2}(v)}\mathbb{E}\left[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v\right]p_{V}(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+O(h^{2})$
$\textstyle=1+O(h^{2})$ (29)
by the definition of $\sigma_{Q}^{2}(v)$. The quantity from Lemma 17 is $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}$. We apply the Lindeberg-Feller Theorem (see for example [20], Thm. 3.4.10.) to the random variables $(W_{i}-\mathbb{E}[W_{i}])/\sqrt{N}$. For this, note that
$\sum_{i=1}^{N}\mathbb{E}\left[(W_{i}-\mathbb{E}[W_{i}])^{2}\right]/N=\mathbb{E}[W_{i}^{2}]-\mathbb{E}[W_{i}]^{2}=1+O(h^{2}).$
Moreover, for all $\delta>0$, using the inequality $|a+b|^{\gamma}\leq 2^{\gamma}(|a|^{\gamma}+|b|^{\gamma})$ for $\gamma>0$ and Jensen’s inequality
$\textstyle\sum_{i=1}^{N}\mathbb{E}\left[\frac{(W_{i}-\mathbb{E}[W_{i}])^{2}}{N}\mathbf{1}\{|W_{i}-\mathbb{E}[W_{i}]|>\sqrt{N}\delta\}\right]$
$\textstyle=\mathbb{E}\left[{(W_{i}-\mathbb{E}[W_{i}])^{2}}\mathbf{1}\{|W_{i}-\mathbb{E}[W_{i}]|>\sqrt{N}\delta\}\right]$
$\textstyle\leq\frac{\mathbb{E}\left[|W_{i}-\mathbb{E}[W_{i}]|^{2+\eta}\right]}{\delta^{\eta}N^{\eta/2}}$
$\textstyle\leq 2^{2+\eta}\frac{\mathbb{E}\left[|W_{i}|^{2+\eta}\right]+|\mathbb{E}[W_{i}]|^{2+\eta}}{\delta^{\eta}N^{\eta/2}}$
$\textstyle\leq 2^{3+\eta}\frac{\mathbb{E}\left[|W_{i}|^{2+\eta}\right]}{\delta^{\eta}N^{\eta/2}}$
$\textstyle=2^{3+\eta}\frac{\mathbb{E}\left[\left|\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)\right|^{2+\eta}\right]}{{\delta^{\eta}N^{\eta/2}h^{\eta/2}\sigma_{Q}(v)^{2+\eta}}\cdot h}.$
This converges to 0 for $N\to\infty$, because $Nh\to\infty$ and because of assertion 12 of Assumption 6. From the Lindeberg-Feller Theorem, we can conclude that
$\frac{1}{\sqrt{N}}\sum_{i=1}^{N}(W_{i}-\mathbb{E}[W_{i}])=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}-\sqrt{N}\mathbb{E}[W_{i}]\to\mathcal{N}(0,1)$
in distribution. By (28) and Assumption 4, it follows that also $\sum_{i=1}^{N}W_{i}/\sqrt{N}\to\mathcal{N}(0,1),$ which concludes to Proof of Lemma 17.
D.1.2 Proof of Lemma 18
Define
$\Delta l_{i}^{k}$ $=l(X_{i})-\hat{l}^{k}(X_{i}),$ $\quad\Delta f_{i}^{k}$ $=f(Z_{i},X_{i})-\hat{f}^{k}(Z_{i},X_{i}),$
$\Delta\phi_{1,i}^{k}$ $=\phi(X_{i})-\hat{\phi}_{1}^{k}(X_{i}),$ $\quad\Delta\phi_{2,i}^{k}$ $=\phi(X_{i})-\hat{\phi}_{2}^{k}(X_{i})$
Under $H_{0}:\beta(v)=\beta_{0}(v)$ we can write (using that $Y_{i}-l(X_{i})-\beta(v)(D_{i}-\phi(X_{i}))=\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))$ in the last step)
$\textstyle\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\sigma_{Q}(v)}$
$=\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))\cdot$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$=\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-l(X_{i})-\beta(v)(D_{i}-\phi(X_{i}))+\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\cdot$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad\cdot\left(f(Z_{i},X_{i})-\phi(X_{i})-\Delta f_{i}^{k}+\Delta\phi_{2,i}^{k}\right)K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle=\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{i=1}^{N}\left(Y_{i}-l(X_{i})-\beta(v)(D_{i}-\phi(X_{i}))\right)\cdot$
$\textstyle\quad\qquad\qquad\qquad\quad\cdot(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle+\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\Bigl[\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)\Bigr]K\left(\frac{V_{i}-v}{h}\right).$ (30)
Hence, to arrive at Lemma 18, we need to show that for $k=1,\ldots,K$,
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (31)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (32)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (33)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta\phi_{i,2}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (34)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (35)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{i,1}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (36)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (37)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (38)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta\phi_{1,i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (39)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta\phi_{1,i}^{k}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1).$ (40)
Recall from Section 2.4 that by conditioning on $I_{k}^{c}$, we mean conditioning on $\{X_{i},Y_{i},Z_{i},D_{i}\}_{i\in I_{k}^{c}}$. For (31), observe that
$\textstyle\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c},X_{I_{k}},Z_{I_{k}}\right]$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\sum_{j\in I_{k}}\mathbb{E}\left[\epsilon_{i}\epsilon_{j}\Delta f_{i}^{k}\Delta f_{j}^{k}K\left(\frac{V_{i}-v}{h}\right)K\left(\frac{V_{j}-v}{h}\right)|I_{k}^{c},X_{i},X_{j},Z_{i},Z_{j}\right]$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\sum_{j\in I_{k}}\mathbb{E}\left[\epsilon_{i}\epsilon_{j}|X_{i},X_{j},Z_{i},Z_{j}\right]\Delta f_{i}^{k}\Delta f_{j}^{k}K\left(\frac{V_{i}-v}{h}\right)K\left(\frac{V_{j}-v}{h}\right)$
$\textstyle=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\left(\Delta f_{i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ (41)
$\textstyle\leq\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\left(\Delta f_{i}^{k}\right)^{2}.$
It follows that
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c}\right]$
$\leq\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\left(\Delta f_{i}^{k}\right)^{2}|I_{k}^{c}\right]$
$=\|K\|_{\infty}^{2}\|\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\|_{L_{\infty}}^{2}\frac{\|f-\hat{f}^{k}\|_{L_{2}}^{2}}{h\sigma_{Q}^{2}(v)}$
By assertions 2 and 4 of Assumption 5, the last expression is $o_{P}(1)$. By Lemma 24 from Section D.5, we obtain (31). In exactly the same way, one can also show (32).
For (33), we use the Cauchy-Schwarz inequality in the second step and assertion 8 of Assumption 6 with Lemma 23 in the fourth step to obtain
$\mathbb{E}\left[\left|\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\left|(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\left|(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right||I_{k}^{c}\right]^{1/2}\cdot$
$\qquad\qquad\qquad\qquad\quad\cdot\mathbb{E}\left[\left(\Delta f_{i}^{k}\right)^{2}|I_{k}^{c}\right]^{1/2}$
$=\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[(\beta(V_{i})-\beta(v))^{2}\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}|V_{i}\right]K\left(\frac{V_{i}-v}{h}\right)^{2}\right]^{1/2}\cdot$
$\qquad\qquad\qquad\cdot\|f-\hat{f}^{k}\|_{L_{2}}$
$=\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(v)}O(h^{3/2})\|f-\hat{f}^{k}\|_{L_{2}}$
$=\frac{\|f-\hat{f}^{k}\|_{L_{2}}}{\sigma_{Q}(v)}O(n^{1/2}h),$
which is $o_{P}(1)$ by assertion 1 of Assumption 5. By Lemma 24, (33) follows. In exactly the same way, we can also show (34).
For (35), we can argue similarly to (31). Concretely,
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c},X_{I_{k}}\right]$
$=\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\left(\Delta l_{i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ (42)
and hence using Hölder’s inequality with $1/p+1/q=1$
$\mathbb{E}\left[\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}|I_{k}^{c}\right]$
$=\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\left(\Delta l_{i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}|I_{k}^{c}\right]$
$\leq\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]^{p}K\left(\frac{V_{i}-v}{h}\right)^{2p}\right]^{1/p}\mathbb{E}\left[|\Delta l_{i}^{k}|^{2q}|I_{k}^{c}\right]^{1/q}$
$\leq\frac{\|K\|_{\infty}^{2}}{h\sigma_{Q}^{2}(v)}\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p}}\|l-\hat{l}^{k}\|_{L_{2q}}^{2}$
If we set $p=p_{1}$ and $q=q_{1}$, this is $o_{P}(1)$ using assertion 5 of Assumption 5, and (35) follows using Lemma 24. In exactly the same way, (36) follows.
For (37), we use the Cauchy-Schwarz inequality to obtain
$\mathbb{E}\left[\left|\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right||I_{k}^{c}\right]$
$\leq\frac{\sqrt{n}}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[\left|\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}\right|K\left(\frac{V_{i}-v}{h}\right)|I_{k}^{c}\right]$
$\leq\frac{\sqrt{n}\|K\|_{\infty}}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[\left(\Delta l_{i}^{k}\right)^{2}|I_{k}\right]^{1/2}\mathbb{E}\left[\left(\Delta\phi_{2,i}^{k}\right)^{2}|I_{k}^{c}\right]^{1/2}$
$=\frac{\sqrt{n}\|K\|_{\infty}}{\sqrt{h}\sigma_{Q}(v)}\|l-\hat{l}^{k}\|_{L_{2}}\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$
$=o_{P}(1)$
by assertion 3 of Assumption 5. Hence, (37) follows using Lemma 24. In exactly the same way, also (38), (39) and (40) follow. This concludes the proof of Lemma 18.
D.2 Proof or Theorem 12
By Theorem 11, $h\hat{Q}(\beta_{0}(v),v)/\sigma_{Q}(v)=O_{P}\left(h/\sqrt{Nh}\right)=o_{P}(1)$. Hence it is enough to show the following two Lemmas.
Lemma 19.
Under the conditions of Theorem 12,
$\textstyle\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\textstyle=1+o_{P}(1).$
Lemma 20.
Under the conditions of Theorem 12,
$\textstyle\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta_{0}(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))^{2}\cdot$
$\textstyle\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\textstyle-\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))^{2}\cdot$
$\textstyle\qquad\qquad\qquad\quad\cdot(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\textstyle=o_{P}(1).$
D.2.1 Proof of Lemma 19
As in the proof of Lemma 17, define
$W_{i}=\frac{1}{\sqrt{h}\sigma_{Q}(v)}(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right).$ (43)
We need to show $\frac{1}{N}\sum_{i=1}^{N}W_{i}^{2}-1=o_{P}(1)$. By (29), $\mathbb{E}[W_{i}^{2}]=1+O(h^{2})=1+o_{P}(1)$. Hence, it suffices to show that $\frac{1}{N}\sum_{i=1}^{N}\left(W_{i}^{2}-\mathbb{E}[W_{i}^{2}]\right)=o_{P}(1).$ By Chebyshev’s Inequality, for all $\delta>0$
$\mathbb{P}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\left(W_{i}^{2}-\mathbb{E}[W_{i}^{2}]\right)\right|>\delta\right)\leq\frac{\mathrm{Var}(W_{i}^{2})}{N\delta}\leq\frac{\mathbb{E}[W_{i}^{4}]}{N\delta}=o(1)$
by assertion 13 of Assumption 6. This concludes the proof of Lemma 19.
D.2.2 Proof of Lemma 20
Define $W_{i}$ as in (43) and for $i\in I_{k}$, $k=1,\ldots,K$
$\Gamma_{i}^{k}$ $=-W_{i}+\frac{1}{\sqrt{h}\sigma_{Q}(v)}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))\cdot$
$\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right).$
Using this notation, we need to show that
$\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(\left(\Gamma_{i}^{k}+W_{i}\right)^{2}-W_{i}^{2}\right)=o_{P}(1).$
By the Cauchy-Schwarz inequality,
$\left|\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(\left(\Gamma_{i}^{k}+W_{i}\right)^{2}-W_{i}^{2}\right)\right|=\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left|\left(\Gamma_{i}^{k}\right)^{2}+2\Gamma_{i}^{k}W_{i}\right|$
$\leq\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(\Gamma_{i}^{k}\right)^{2}+2\sqrt{\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(\Gamma_{i}^{k}\right)^{2}}\sqrt{\frac{1}{N}\sum_{i=1}^{N}W_{i}^{2}}.$
From Lemma 19, we know that $\frac{1}{N}\sum_{i=1}^{N}W_{i}^{2}=1+o_{P}(1)$ and hence it remains to show that $\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(\Gamma_{i}^{k}\right)^{2}=o_{P}(1)$. We use the same decomposition as in (30) to see that
$\Gamma_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}\Bigl[\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)$
$\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)$
$\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)\Bigr]K\left(\frac{V_{i}-v}{h}\right).$
Hence, to arrive at Lemma 20, it suffices to show that for $k=1,\ldots,K$,
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (44)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (45)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (46)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (47)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta l_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (48)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (49)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (50)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (51)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta\phi_{1,i}^{k})^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1)$ (52)
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta\phi_{1,i}^{k})^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=o_{P}(1).$ (53)
For (44), observe that
$\mathbb{E}\left[\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\epsilon_{i}^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}|I_{k}^{c},Z_{I_{k}},X_{I_{k}}\right]$
$=\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\mathbb{E}\left[\epsilon_{i}^{2}|Z_{i},X_{i}\right](\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2},$
which is the same expression as (41). Hence, (44) follows as before. The same argument can be applied for (45). From the proof of (33), we know that
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|=o_{P}(1).$
Hence, (46) follows, since
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\leq\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|\right)^{2}.$
In the same way, also (47) follows. For (48), we argue as for (44) by noting that the conditional expectation of the expression given $(I_{k}^{c},X_{I_{k}})$ is equal to (42) that has been analyzed before. The same reasoning applies to (49). For (50), we have established in the proof of (37) that
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|=o_{P}(1).$
Hence, (50) follows by noting that
$\frac{1}{{nh}\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta l_{i}^{k})^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$\leq\left(\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\left|\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|\right)^{2}.$
In the same way, also (51), (52) and (53) follow.
D.3 Proof of Theorem 8
We can write
$\textstyle\frac{\sqrt{Nh}(\hat{\beta}(v)-\beta(v))}{\sigma(v)}$
$\textstyle=\frac{\frac{1}{\sqrt{Nh}}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)}{\frac{\sigma(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)}$
$\textstyle=\frac{\sqrt{Nh}\hat{Q}(\beta(v),v)/\sigma_{Q}(v)}{\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)}$
From Theorem 11, we know that the numerator converges in distribution to $\mathcal{N}(0,1)$. By Slutsky’s Theorem, it is enough to show that the denominator converges to $1$ in probability. Hence, it is enough to prove the following two lemmas.
Lemma 21.
Under the conditions of Theorem 8,
$\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)-1=o_{P}(1).$
Lemma 22.
Under the conditions of Theorem 8,
$\textstyle\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle-\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$ $\textstyle=o_{P}(1).$
D.3.1 Proof of Lemma 21
Note that by the tower property of conditional expectations,
$\frac{\sigma(v)}{\sigma_{Q}(v)}$ $=\left(\mathbb{E}[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))|V_{i}=v]p_{V}(v)\right)^{-1}$
$=\left(\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]p_{V}(v)\right)^{-1}.$ (54)
By Lemma 23 and assertion 9 of Assumption 6
$\mathbb{E}\left[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)\right]$
$=\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}]K\left(\frac{V_{i}-v}{h}\right)\right]$
$=h\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]p_{V}(v)+\sigma_{Q}^{2}(v)O(h^{3}).$
From assertions 1 and 2 of Assumption 6, it follows that
$U_{i}=\frac{\sigma(v)/\sigma_{Q}(v)}{h}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
satisfy $\mathbb{E}[U_{i}]=1+O(h^{2})=1+o(1)$. To prove Lemma 21, it therefore remains to prove $\frac{1}{N}\sum_{i=1}^{N}(U_{i}-\mathbb{E}[U_{i}])=o_{P}(1)$. For this, using Chebyshev’s inequality, it is enough to show
$\frac{\mathbb{E}[U_{i}^{2}]}{N}=o(1).$ (55)
We calculate $\frac{\mathbb{E}[U_{i}^{2}]}{N}$ using (54), apply Lemma 23 and assertions 10, 3, 2 and 1 of Assumption 6 to get
$\textstyle\frac{\mathbb{E}[U_{i}^{2}]}{N}$ $=\textstyle\frac{\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]}{Nh^{2}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]^{2}p_{V}(v)^{2}}$
$\textstyle=\frac{h\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v\right]p_{V}(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+\sigma_{Q}^{2}(v)O(h^{3})}{Nh^{2}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]^{2}p_{V}(v)^{2}}$
$\textstyle\lesssim\frac{h\sigma_{Q}^{2}(v)+h^{3}\sigma_{Q}^{2}(v)}{Nh^{2}\sigma_{Q}^{4}(v)}$
$\textstyle\lesssim\frac{1}{Nh\sigma_{Q}^{2}(v)}.$
Since by the assumption of Theorem 8, $\sigma_{Q}^{2}(v)\gg(Nh)^{-1}$, (55) follows. This concludes the proof of Lemma 21.
D.3.2 Proof of Lemma 22
Let $\Delta\phi_{1,i}^{k}=\phi(X_{i})-\hat{\phi}_{1}^{k}(X_{i})$, $\Delta\phi_{2,i}^{k}=\phi(X_{i})-\hat{\phi}_{2}^{k}(X_{i})$ and $\Delta f_{i}^{k}=f(X_{i})-\hat{f}^{k}(X_{i})$ as before. Then, the quantity given in Lemma 22 is equal to
$\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\bigl[(D_{i}-\phi(X_{i}))(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k})$
$\quad\quad\quad+\Delta\phi_{1,i}^{k}(f(Z_{i},X_{i})-\phi(X_{i}))$
$\quad\quad\quad+\Delta\phi_{1,i}^{k}(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{i})\bigr]K\left(\frac{V_{i}-v}{h}\right).$
Moreover, we know from (54) and assertions 1 and 2 of Assumption 6 that $\sigma(v)/\sigma_{Q}(v)\asymp\frac{1}{\sigma_{Q}^{2}(v)}$ Hence, it remains to show that for $k=1,\ldots,K$,
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (56)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (57)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{1,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (58)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\Delta f_{i}^{k}\Delta\phi_{1,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (59)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\Delta\phi_{1,i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1).$ (60)
For (56), we use the Cauchy-Schwarz inequality to get
$\left|\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|$
$\leq\frac{1}{\sigma_{Q}^{2}(v)}{\sqrt{\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)}}{\sqrt{\frac{1}{nh}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)}}.$
Hence, to show (56), it is enough to prove that
$\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)$ $=O_{P}(1)$ (61)
$\frac{1}{\sigma_{Q}^{4}(v)nh}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1).$ (62)
(61) can be obtained as in the proof of Lemma 19 using Assertions 11 and 14 of Assumption 6 and Lemma 23. (62) follows similarly to the proof of (37) by noticing that $\frac{\|K\|_{\infty}}{\sigma_{Q}^{4}(v)h}\|f-\hat{f}^{k}\|_{L_{2}}^{2}=o_{P}(1)$ by the assumption in Theorem 8. Hence, we obtain (56). In the same way, one can obtain (57). For (58), it follows from (36) that
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{1,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=\frac{1}{\sigma_{Q}(v)\sqrt{nh}}o_{P}(1).$
By the assumption of Theorem 8, $\sigma_{Q}^{2}(v)\gg(Nh)^{-1}$ and hence, (58) follows. In the same way, (59) and (60) follow from (40) and (39).
D.4 Proof of Theorem 9
We know from Lemma 21 and Lemma 22 that for the denominator of $\hat{\sigma}^{2}(v)$ defined in (24),
$\left(\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)\right)^{2}$
$=1+o_{P}(1).$
Therefore, it remains to show that for the numerator we have
$\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\hat{\beta}(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)^{2}\cdot$
$\qquad\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$=1+o_{P}(1).$
This equality was shown to hold if one replaces $\hat{\beta}(v)$ by $\beta(v)$ in Lemma 19 and Lemma 20. Hence, it is enough to show
$\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\hat{\beta}(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)^{2}\cdot$ (63)
$\qquad\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$-\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)^{2}\cdot$ (64)
$\qquad\qquad\qquad\qquad\quad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$=o_{P}(1).$ (65)
Define
$\Xi_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}(\beta(v)-\hat{\beta}(v))(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\cdot$
$\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)$
$\Theta_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}\left(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)\cdot$
$\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right).$
The expression in (65) is then equal to
$\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left((\Xi_{i}^{k}+\Theta_{i}^{k})^{2}-(\Theta_{i}^{k})^{2}\right).$
By the same argument as in the proof of Lemma 20, it remains to show
$\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}(\Theta_{i}^{k})^{2}$ $=O_{P}(1)$ (66)
$\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in I_{k}}(\Xi_{i}^{k})^{2}$ $=o_{P}(1)$ (67)
(66) holds by Lemma 19 and Lemma 20. For (67), observe that
$\frac{1}{N}\sum_{i=1}^{N}(\Xi_{i}^{k})^{2}$ $=\frac{(\beta(v)-\hat{\beta}(v))^{2}}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))^{2}\cdot$
$\qquad\qquad\qquad\qquad\qquad\quad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}.$
By Theorem 8, we know that $\sqrt{Nh}(\hat{\beta}(v)-\beta(v))/\sigma(v)=O_{P}(1)$. From (54) and assertions 1 and 2 of Assumption 6, we know that $\sigma(v)\asymp 1/\sigma_{Q}(v)$. Since by assumption $\sigma_{Q}(v)\gg(Nh)^{-1/2}$, it follows that $(\hat{\beta}(v)-\beta(v))^{2}=o_{P}(1)$. It remains to prove that
$\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))^{2}\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$=O_{P}(1).$ (68)
As before, we first consider the “oracle” version of (68) with $\hat{\phi}_{1,2}^{k}$ and $\hat{f}^{k}$ replaced by $\phi$ and $f$. For this, note that by assertion 10 of Assumption 6 and Lemma 23,
$\mathbb{E}\left[\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))^{2}\left(f(Z_{i},X_{i})-\phi(X_{i})\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$ $=O(1)$
and by Lemma 25, we conclude that the oracle version of (68) is $O_{P}(1)$. It remains to show that the difference between (68) and its “oracle” version is $O_{P}(1)$ (note that we only need $O_{P}(1)$ here and not $o_{P}(1)$). Similarly to before, this reduces to show for $k=1,\ldots,K$
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (69)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (70)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (71)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}(\Delta\phi_{1,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1)$ (72)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(\Delta\phi_{1,i}^{k})^{2}(\Delta\phi_{2,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$ $=O_{P}(1).$ (73)
For (69), using the Cauchy-Schwarz inequality, it is enough to have $\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{4}K\left(\frac{V_{i}-v}{h}\right)^{2}=O_{P}(1)$ and $\frac{1}{nh\sigma_{Q}^{4}(v)}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{4}K\left(\frac{V_{i}-v}{h}\right)^{2}=O_{P}(1)$. The expectation of the first term is bounded by assertion 14 of Assumption 6 and hence the term is $O_{P}(1)$ by Lemma 25. The second follows by taking the expectation conditional on $I_{k}^{c}$ since $\|f-\hat{f}\|_{L_{4}}/\sigma_{Q}(v)=O_{P}\left(h^{1/4}\right)$ by the assumption of Theorem 9. Hence, (69) follows and (70) follows in exactly the same way. Finally, (71), (72) and (73) follow from (49), (53) and (52). This concludes the proof of (68) and hence also the proof of Theorem 9.
D.5 Some Lemmas
Lemma 23.
Assume that the kernel $K$ satisfies Assumption 3. Moreover, assume that the function $\rho:\mathbb{R}\to\mathbb{R}$ defined by $\rho(x)=v(x)\cdot p_{V}(x)$ for some function $v(\cdot)$ is two times differentiable for $x$ in the support of $V_{i}$. Then,
  1. 1.
    $\mathbb{E}\left[K\left(\frac{V_{i}-v}{h}\right)v(V_{i})\right]=h\rho(v)+h^{3}\frac{\rho^{\prime\prime}(v)}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x+R(v,h)$
    with
    $|R(v,h)|\leq h^{3}\Biggl(\|\rho^{\prime\prime}\|_{\infty}\int_{|x|>h^{-1/2}}K(x)x^{2}\mathrm{d}x$
    $\qquad\qquad\qquad\quad+\frac{1}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x\cdot\sup_{|s-v|\leq h^{1/2}}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\Biggr).$
    In particular, if $\rho^{\prime\prime}$ is continuous at $v$, then $R(v,h)=o(h^{3})$. If $|\rho^{\prime\prime}|$ bounded but not continuous at $v$, we still have $R(v,h)=O(h^{3})$.
  2. 2.
    $\mathbb{E}\left[K\left(\frac{V_{i}-v}{h}\right)^{2}v(V_{i})\right]$
    $=h\rho(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+h^{3}\frac{\rho^{\prime\prime}(v)}{2}\int_{-\infty}^{\infty}K(x)^{2}x^{2}\mathrm{d}x+\bar{R}(v,h)$
    with
    $|\bar{R}(a,h)|\leq h^{3}\Biggl(\|\rho^{\prime\prime}\|_{\infty}\int_{|x|>h^{-1/2}}K(x)^{2}x^{2}\mathrm{d}x$
    $\qquad\qquad\qquad\quad+\frac{1}{2}\int_{-\infty}^{\infty}K(x)^{2}x^{2}\mathrm{d}x\cdot\sup_{|s-v|\leq h^{1/2}}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\Biggr).$
    In particular, if $\rho^{\prime\prime}$ is continuous at $v$, then $\bar{R}(v,h)=o(h^{3})$. If $|\rho^{\prime\prime}|$ bounded but not continuous at $v$, we still have $\bar{R}(v,h)=O(h^{3})$.
Proof.
We only prove the first statement. The second statement follows by replacing $K$ with $K^{2}$ in the proof. From Taylor’s theorem, we can write for some $\xi_{x}\in[v,x]$ (or in $[x,v]$)
$\rho(x)=\rho(v)+\rho^{\prime}(v)(x-v)+\rho^{\prime\prime}(v)\frac{(x-v)^{2}}{2}+(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v))\frac{(x-v)^{2}}{2}.$
Hence, using that $K(\cdot)$ is symmetric,
$\mathbb{E}\left[K\left(\frac{V_{i}-v}{h}\right)v(V_{i})\right]$ $=\int K\left(\frac{x-v}{h}\right)v(x)p_{V}(x)\mathrm{d}x$
$=\int K\left(\frac{x-v}{h}\right)\rho(x)\mathrm{d}x$
$=\rho(v)\int K\left(\frac{x-v}{h}\right)\mathrm{d}x$
$+\frac{\rho^{\prime\prime}(v)}{2}\int K\left(\frac{x-v}{h}\right)(x-v)^{2}\mathrm{d}x+R(v,h)$
$=h\rho(v)+h^{3}\frac{\rho^{\prime\prime}(v)}{2}\int K(x)x^{2}\mathrm{d}x+R(v,h)$
with
$R(v,h)=\int K\left(\frac{x-v}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v)\right)\frac{(x-v)^{2}}{2}\mathrm{d}x.$
To conclude, note that for any $t>0$,
$|R(v,h)|$
$\leq|\int_{|x-v|>th}K\left(\frac{x-v}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v)\right)\frac{(x-v)^{2}}{2}\mathrm{d}x|$
$+|\int_{|x-v|\leq th}K\left(\frac{x-v}{h}\right)\left(\rho^{\prime\prime}(\xi_{x})-\rho^{\prime\prime}(v)\right)\frac{(x-v)^{2}}{2}\mathrm{d}x|$
$\leq\int_{|x-v|>th}K\left(\frac{x-v}{h}\right)2\|\rho^{\prime\prime}\|_{\infty}\frac{(x-v)^{2}}{2}\mathrm{d}x$
$+\int_{|x-v|\leq th}K\left(\frac{x-v}{h}\right)\sup_{|s-a|\leq ht}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\frac{(x-v)^{2}}{2}\mathrm{d}x$
$\leq h^{3}\left(\|\rho^{\prime\prime}\|_{\infty}\int_{|x|>t}K(x)x^{2}\mathrm{d}x+\frac{1}{2}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x\cdot\sup_{|s-v|\leq th}|\rho^{\prime\prime}(s)-\rho^{\prime\prime}(v)|\right).$
Choosing $t=h^{-1/2}$ gives the result. ∎
Lemma 24.
Let $(X_{n})_{n\in\mathbb{N}}$ be a sequence of integrable random variables and $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ be a sequence of sigma-algebras. If $\mathbb{E}[|X_{n}||\mathcal{F}_{n}]=o_{P}(1)$, then we have that $X_{n}=o_{P}(1)$.
Proof.
Let $\epsilon>0$. By Markov’s inequality,
$\mathbb{P}(|X_{n}|\geq\epsilon)=\mathbb{P}(|X_{n}|\land\epsilon\geq\epsilon)\leq\mathbb{E}\left[\frac{\mathbb{E}[|X_{n}|\land\epsilon|\mathcal{F}_{n}]}{\epsilon}\right]\leq\mathbb{E}\left[\frac{\mathbb{E}[|X_{n}||\mathcal{F}_{n}]}{\epsilon}\land 1\right]$
Since $\mathbb{E}[|X_{n}||\mathcal{F}_{n}]=o_{P}(1)$, for each $\delta\in(0,1)$, there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, we have $\mathbb{P}(\mathbb{E}[|X_{n}||\mathcal{F}_{n}]\geq\epsilon\delta/2]\leq\delta/2$. It follows that for all $n\geq n_{0}$,
$\mathbb{P}(|X_{n}|\geq\epsilon)\leq\frac{\epsilon\delta/2}{\epsilon}\mathbb{P}\left(\mathbb{E}[|X_{n}||\mathcal{F}_{n}]\leq\epsilon\delta/2\right)+\mathbb{P}\left(\mathbb{E}[|X_{n}||\mathcal{F}_{n}]\geq\epsilon\delta/2\right)\leq\delta.$
Since $\epsilon$ and $\delta$ were arbitrary, it follows that $X_{n}=o_{P}(1)$. ∎
Lemma 25.
Let $(X_{n})_{n\in\mathbb{N}}$ be a sequence of integrable random variables and $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ be a sequence of sigma-algebras. If $\mathbb{E}[|X_{n}||\mathcal{F}_{n}]=O_{P}(1)$, then we have that $X_{n}=O_{P}(1)$.
Proof.
Let $\epsilon\in(0,1)$. Since $\mathbb{E}[|X_{n}||\mathcal{F}_{n}]=O_{P}(1)$, there exists $M>0$ and $n_{0}\in\mathbb{N}$, such that for all $n\geq n_{0}$
$\mathbb{P}(\mathbb{E}[|X_{n}||\mathcal{F}_{n}]>\epsilon M/2)\leq\epsilon/2.$
By Markov’s inequality, for all $n\geq n_{0}$,
$\mathbb{P}(|X_{n}|\geq M)$ $=\mathbb{P}(|X_{n}|\land M\geq M)$
$\leq\mathbb{E}\left[\frac{\mathbb{E}[|X_{n}|\land M|\mathcal{F}_{n}]}{M}\right]$
$\leq\mathbb{E}\left[\frac{\mathbb{E}[|X_{n}||\mathcal{F}_{n}]}{M}\land 1\right]$
$\leq\mathbb{P}\left(\frac{\mathbb{E}[|X_{n}||\mathcal{F}_{n}]}{M}>\epsilon/2\right)+\epsilon/2\leq\epsilon.$
Since $\epsilon\in(0,1)$ was arbitrary, $X_{n}=O_{P}(1)$. ∎
Appendix E Further Simulations
E.1 Varying IV Strength
We consider the same simulation setup (26) as in Section 4.2 with two differences:
  1. 1.
    We introduce an additional parameter $s$ (for the IV strength) to the function $f(z,x)$, namely
    (Z lin.)
    $f(z,x)=-\sin(x)+s\cdot z$,
    (Z nonlin.)
    $f(z,x)=-\sin(x)+s\cdot(\cos(z)+0.2z)$
    If the strength $s=0$, the treatment $D_{i}$ is not related to the instrument $Z_{i}$. If $s$ increases, the strength of the instrument $Z_{i}$ increases.
  2. 2.
    We replace the definitions of $\delta_{i}$ and $\epsilon_{i}$ by
    $\delta_{i}$ $\leftarrow 0.7H_{i}+0.1E_{\delta,i},$
    $\epsilon_{i}$ $\leftarrow 0.7H_{i}+0.1E_{\epsilon,i}$
    to make them more correlated, i.e., increase the amount of endogeneity.
We fix $N=500$ and vary the strength $s\in\{0,0.1,\ldots,1\}$. Again, we simulate $n_{\text{sim}}=1000$ datasets and use generalized additive models (gam) for the nuisance functions, the Epanechnikov kernel, and choose the bandwidth as the normal reference rule times $N^{1/5}/N^{1/7}$. The corresponding plots for the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.) can be found in Figure 6 and the plots for the settings (het.)/(Z lin.) and (het.)/(Z nonlin.) can be found in Figure 7.
Figure 6: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom) with varying IV strength $s$. Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Figure 7: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom) with varying IV strength $s$. Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
In both figures, we see that the mean squared error tends to decrease with increasing IV strength. Moreover, for small to moderately small IV strengths, the standard confidence intervals suffer from severe undercoverage and for the (Z nonlin.) settings, the undercoverage is worse for the methods based on (linearIV). In contrast, the robust confidence sets exhibit empirical coverage close to the nominal level of 0.95 regardless of the IV strength.
E.2 Multidimensional $X$
We keep the framework of Section 4.2 but use 5-dimensional covariates $X_{i}$ and use boosted regression trees using the R package xgboost [12, 13] for the nuisance function estimations. We slightly alter the simulation setup (26) and generate data according to
$X_{i}$ $\sim_{i.i.d.}\mathcal{N}_{5}(0,\Sigma),\,\Sigma_{i,j}=0.5+0.5\cdot 1\{i=j\},$
$H_{i},E_{Z,i},E_{\delta,i},E_{\epsilon,i}$ $\sim_{i.i.d.}\mathcal{N}(0,1),$
$Z_{i}$ $\leftarrow 0.5X_{i,1}-0.5X_{i,2}+E_{Z,i},$
$\delta_{i}$ $\leftarrow 0.7H_{i}+0.7E_{\delta,i},$ (74)
$\epsilon_{i}$ $\leftarrow\mathrm{sign}(H_{i})-0.5+0.5E_{\epsilon,i},$
$D_{i}$ $\leftarrow f(Z_{i},X_{i})+\delta_{i},$
$V_{i}$ $\leftarrow X_{i,1},$
$Y_{i}$ $\leftarrow\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i}.$
Note that the components of $X_{i}$ all have variance 11 and are correlated. Moreover, we change the function $g$ to $g(x)=\tanh(x_{1})-x_{3}$ and change the settings (Z lin.) and (Z nonlin.) to
(Z lin.)
$f(z,x)=-\sin(x_{1})+x_{2}+z$,
(Z nonlin.)
$f(z,x)=-\sin(x_{1})+x_{2}+\cos(z)+0.2z$.
The remaining simulation setup stays the same as in Section 4.2. Again, we first consider the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.) and use the methods (hom. T.E. mlIV) and (hom. T.E. linearIV) as well as the methods (het. T.E. mlIV) and (het. T.E. linearIV) for $\hat{\beta}(0)$ and $\hat{\beta}(1.5)$. The results are shown in Figure 8.
Figure 8: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom) for 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$)and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
The results are very similar to Section 4.2: in the (Z lin.) settings, the estimators based on (linearIV) exhibit almost identical results to the estimators based on (mlIV), whereas in the (Z nonlin.) settings, the estimators based on (mlIV) have significantly smaller MSEs, while still exhibiting good empirical coverage of both standard and robust confidence intervals.
Secondly, we consider the settings (het.)/(Z lin.) and (het.)/(Z nonlin.) and use the methods (het. T.E. mlIV) and (het. T.E. linearIV) for $\hat{\beta}(0)$ and $\hat{\beta}(1.5)$. The results are shown in Figure 9. Again, the results are comparable to those from Section 4.2.
Figure 9: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom) for 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Finally, let us consider the influence of the IV strength in the setting with 5-dimensional $X_{i}$. As in Section E.1, we include a parameter $s$ corresponding to different IV strengths and modify $\delta_{i}$ and $\epsilon_{i}$ to increase their correlation. Otherwise, we keep the setup from (74). The resulting plots for the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.) can be found in Figure 10 and the plots for the settings (het.)/(Z lin.) and (het.)/(Z nonlin.) can be found in Figure 11.
Figure 10: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom) with varying IV strength $s$, 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Figure 11: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom) with varying IV strength $s$, 5-dimensional $X_{i}$ and nuisance function estimation using xgboost. Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
Whereas for the (hom.) settings, there is still a considerable (though less pronounced) undercoverage of the standard confidence intervals for small IV strength, this effect seems to be almost gone for the (het.) settings.